### v1

# The twisted coarse Baum–Connes conjecture and relative hyperbolic groups

Abstract
In this paper, we introduce a notion of stable coarse algebras for metric spaces with bounded geometry, and formulate the twisted coarse Baum–Connes conjecture with respect to stable coarse algebras. We prove permanence properties of this conjecture under coarse equivalences, unions and subspaces.
As an application, we study higher index theory for a group $G$ that is hyperbolic relative to a finite family of subgroups $\{H_{1},H_{2},\dots,H_{N}\}$. We prove that $G$ satisfies the twisted coarse Baum–Connes conjecture with respect to any stable coarse algebra if and only if each subgroup $H_{i}$ does.
1 Introduction
Let $X$ be a metric space with bounded geometry. The coarse Baum-Connes conjecture claims that a certain coarse assembly map
$\mu:KX_{*}(X)\to K_{*}(C^{*}(X))$
is an isomorphism, where the left side is the coarse $K$-homology and the right side is the $K$-theory of the Roe algebra of $X$. This conjecture has many important applications in geometry and topology, including the Gromov–Lawson conjecture on the existence of Riemannian metrics with positive scalar curvature, and the Novikov conjecture on the homotopy invariance of higher signature on oriented closed manifolds. It has been verified for a large class of metric spaces, for example, metric spaces which admit a coarse embedding into Hilbert space [Yu00], and some metric spaces which are not coarsely embeddable into Hilbert space, including relative expanders [DWY23].
To study the coarse Baum–Connes conjecture for the extension $1\to N\to G\to G/N\to 1$ with $N$ and $G/N$ coarsely embeddable into Hilbert space (cf. [AT19]), the first author and L. Guo [DG24] developed a twisted coarse Baum-Connes conjecture with coefficients. In their work, they introduced the notion of a coarse algebra. Roughly speaking, for a metric space $X$ with bounded geometry, a $C^{*}$-algebra $\mathcal{A}$ and the algebra of all compact operators $\mathcal{K}$ on a separable infinite dimensional Hilbert space $\mathcal{H}$, a coarse algebra $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$ is a $C^{*}$-subalgebra of $\ell^{\infty}(X,\mathcal{A}\otimes\mathcal{K})$ which is invariant under the pullback by partial translations. This is a coarse geometric counter-part of a coefficient of a crossed product $C^{*}$-algebra. Then they formulated the twisted Roe algebra $C^{*}(X;\Gamma(X,\mathcal{A}\otimes\mathcal{K}))$ associated with a coarse algebra $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$. Using this framework, they verified the coarse Baum-Connes conjecture holds for a metric space with a coarse fibration for which each fiber and the base space are coarsely embeddable into Hilbert space, and this applies to the extension of coarse embeddable groups.
Based on the aforementioned framework of the twisted coarse Baum–Connes conjecture, we introduce a notion of stable coarse algebra for a metric space with bounded geometry and define a refined version of twisted coarse Baum–Connes conjecture. This refined version of twisted coarse Baum–Connes conjecture has better permanence properties, for instance, inheritance to subspaces and coarse invariance. We say that a coarse algebra $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$ is stable if it is invariant under the multiplication in $B(\mathcal{H})$-direction and under forming a matrix (Definition 2.1). We can generalize the twisted Roe algebra to non-discrete metric spaces with bounded geometry. Let $M$ be a metric space and $X\subset M$ a net in $M$. Given a stable coarse $X$-algebra $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$, we can define the $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$-twisted Roe algebra $C^{*}(M;\Gamma(X,\mathcal{A}\otimes\mathcal{K}))$.
In this paper, we study twisted Roe algebras $C^{*}(M;\Gamma(X,\mathcal{A}\otimes\mathcal{K}))$ associated with a stable coarse algebra $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$.
Theorem 1.1.
Let $M$ and $N$ be metric spaces with bounded geometry with fixed nets $X\subset M$ and $Y\subset N$, respectively, and let $f:M\to N$ be a coarse equivalence. For any stable coarse $Y$-algebra $\Gamma(Y,\mathcal{A}\otimes\mathcal{K})$, there exists a stable coarse $X$-algebra $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$ such that
$C^{*}(M;\Gamma(X,\mathcal{A}\otimes\mathcal{K}))\cong C^{*}(N;\Gamma(Y,\mathcal{A}\otimes\mathcal{K})).$
Let $X$ be a discrete metric space with bounded geometry, and $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$ a coarse $X$-algebra. For each Rips complex $P_{s}(X)$, we can define the $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$-twisted localization algebra, denoted by $C^{*}_{L}(P_{s}(X),\Gamma(X,\mathcal{A}\otimes\mathcal{K}))$. This is a $C^{*}$-subalgebra of the localization algebra with coefficients in $\mathcal{A}$. There is a *-homomorphism
$e_{*}:K_{*}(C_{L}^{*}(P_{s}(X);\Gamma(X,\mathcal{A}\otimes\mathcal{K})))\to K_{*}(C^{*}(P_{s}(X);\Gamma(X,\mathcal{A}\otimes\mathcal{K})))\cong K_{*}(C^{*}(X;\Gamma(X,\mathcal{A}\otimes\mathcal{K})))$
induced by the evaluation. Passing to infinity, we obtain the $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$-twisted coarse assembly map
[…]
The twisted coarse Baum–Connes conjecture [DG24] claims that the twisted coarse assembly map is an isomorphism with respect to any coarse $X$-algebra $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$. We study when it is isomorphism for any stable coarse algebras.
The idea of the twisted coarse Baum–Connes conjecture originated from G. Yu’s proof of coarse Baum–Connes conjecture for metric spaces with bounded geometry admitting a coarse embedding into Hilbert space [Yu00].
The twisted coarse Baum–Connes conjecture captures more information about subspaces. We prove the following coarse analogue of the permanence property of Baum–Connes conjecture with coefficients under taking subgroups [CE01].
Theorem 1.2.
Let $X$ be a metric space with bounded geometry and let $Y\subset X$ be a subspace. If $X$ satisfies the twisted coarse Baum–Connes conjecture with respect to any stable coarse algebras, then so does $Y$.
Another aim of this paper is to apply our new framework to study the higher index theory for relatively hyperbolic groups. The concept of relative hyperbolicity has been extensively studied over the past two decades. For a group $G$ that is hyperbolic relative to a finite collection of peripheral subgroups $\{H_{1},\ldots,H_{N}\}$, D. Osin [Osi05] proved that $G$ has finite asymptotic dimension if and only if each subgroup $H_{i}$ does. N. Ozawa [Oza06] established that $G$ has Yu’s Property A if and only if each $H_{i}$ does. Furthermore, M. Dadarlat and E. Guentner [DG07] showed that $G$ admits a coarse embedding into Hilbert space if and only if each $H_{i}$ does.
A natural question is whether the coarse Baum–Connes conjecture holds for the relative hyperbolic group when each subgroup $H_{i}$ satisfies the conjecture. Fukaya and Oguni [FO12] proved that if $G$ is a finitely generated group that is hyperbolic relative to a finite family of subgroups $\{H_{1},\cdots,H_{N}\}$ such that
  1. (1)
    each $H_{i}$ satisfies the coarse Baum–Connes conjecture, and
  2. (2)
    each $H_{i}$ admits a finite $H_{i}$-simplicial complex $\underline{E}H_{i}$ which is a universal space for proper $H_{i}$-actions,
then $G$ satisfies the coarse Baum–Connes conjecture.
In our setting of twisted coarse Baum–Connes conjecture with respect to stable coarse algebras, we are able to remove the second assumption mentioned above. Namely, we obtain the following:
Theorem 1.3.
Let $G$ be a finitely generated group which is hyperbolic relative to a finite family of subgroups $\{H_{1},\cdots,H_{N}\}$. Then $G$ satisfies the twisted coarse Baum–Connes conjecture with respect to any stable coarse algebras if and only if each $H_{i}$ does.
This paper is organized as follows. In section 2, we recall basic concepts of coarse algebras and the twisted coarse Baum–Connes conjecture. Then prove permanence properties of the twisted coarse Baum–Connes conjecture with respect to any stable coarse algebras. In Section 3, we recall controlled $K$-theory and computational techniques. In Section 4, we clarify what we need to prove to complete the proof of Theorem 1.3 using the language of decomposition complexity. In section 5, we complete the proof of Theorem 1.3.
2 The twisted coarse Baum–Connes conjecture with respect to stable coarse algebras
In this section, we review the background on the coarse Baum–Connes conjecture together with its variants. More precisely, we will focus on a version of twisted coarse Baum-Connes conjecture with respect to certain stable coarse algebras. This is a new framework to study the coarse geometry of metric spaces with bounded geometry. Compared with the usual coarse Baum-Connes conjecture, the new framework captures more information of the large scale geometry of the considered space.
2.1 Twisted Roe algebras
In this subsection, we recall the definition of coarse algebras and use them to define the twisted Roe algebra for a metric space. The concept of coarse algebra was introduced by the first-named author and L. Guo in [DG24], and it was used to study the coarse Baum–Connes conjecture for a coarse fibration whose base space and fiber admit a coarse embedding into Hilbert space. We introduce the notion of stable coarse algebras and prove that the corresponding twisted Roe algebras are invariant under coarse equivalence.
Let $X$ be a discrete metric space, and $\mathcal{A}$ a $C^{*}$-algebra. Let $\mathcal{K}$ be the algebra of all compact operators on an infinite-dimensional, separable Hilbert space $\mathcal{H}$ and $\mathcal{A}\otimes\mathcal{K}$ the tensor product. Denoted by $\ell^{\infty}(X,\mathcal{A}\otimes\mathcal{K})$ the $C^{*}$-algebra of all bounded, $(\mathcal{A}\otimes\mathcal{K})$-valued functions on $X$. We fix a *-isomorphism $\psi:\mathcal{K}\otimes\mathcal{K}\to\mathcal{K}$, and also the induced isomorphism ${\rm id}_{\mathcal{A}}\otimes\psi:\mathcal{A}\otimes\mathcal{K}\otimes\mathcal{K}\to\mathcal{A}\otimes\mathcal{K}$ is still denoted by $\psi$.
Let $D,R\subset X$ be subsets of $X$. A bijective map $v:D\to R$ is said to be a partial translation if we have
$\sup_{x\in D}d(x,v(x))<\infty.$
For every function $f\in\ell^{\infty}(X,\mathcal{A}\otimes\mathcal{K})$ and any partial translation $v:D\to R$, we define
$v^{*}f(x)=\begin{cases}f(v(x)),&\mbox{~if~}x\in D\\ 0,&\mbox{otherwise.}\end{cases}$
Now, we are ready to define coarse algebras and stable coarse algebras.
Definition 2.1.
Let $X$ be a discrete metric space and $\mathcal{A}$ a $C^{*}$-algebras. We say that a $C^{*}$-subalgebra $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$ of $\ell^{\infty}(X,\mathcal{A}\otimes\mathcal{K})$ is a coarse $X$-algebra if it satisfies the following conditions:
  1. (1)
    For each $x\in X$, we have that
    $\overline{\{f(x):f\in\Gamma(X,\mathcal{A}\otimes\mathcal{K})\}}=\mathcal{A}\otimes\mathcal{K};$
  2. (2)
    For each $f\in\Gamma(X,\mathcal{A}\otimes\mathcal{K})$ and each partial translation $v:D\to R$, we have that $v^{*}f\in\Gamma(X,\mathcal{A}\otimes\mathcal{K})$.
In addition, a coarse $X$-algebra $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$ is said to be stable if it satisfies the following conditions:
  1. (3)
    For any two bounded sequences $a,b\in\ell^{\infty}(X,B(\mathcal{H}))$, and $f\in\ell^{\infty}(X;\mathcal{A}\otimes\mathcal{K})$, we define a new function $f^{\prime}\in\ell^{\infty}(X;\mathcal{A}\otimes\mathcal{K})$ by
    $f^{\prime}(x)=\left(\text{id}_{\mathcal{A}}\otimes a(x)\right)f(x)\left(\text{id}_{\mathcal{A}}\otimes b(x)\right).$
    Then $f\in\Gamma(X,\mathcal{A}\otimes\mathcal{K})$ implies $f^{\prime}\in\Gamma(X,\mathcal{A}\otimes\mathcal{K})$ for any two bounded sequences $a,b\in\ell^{\infty}(X,B(\mathcal{H}))$;
  2. (4)
    For every $f\in\Gamma(X,\mathcal{A}\otimes\mathcal{K})$ and $k\in\mathcal{K}$, the function $x\mapsto\psi(f(x)\otimes k)$ is an element in $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$.
We remark here that Conditions $(3)$ and $(4)$ are used to guarantee the coarse invariance of twisted Roe algebras.
Examples 2.2.
Let $X$ be a discrete metric space and $\mathcal{A}$ a $C^{*}$-algebra.
  1. (1)
    The algebra $\ell^{\infty}(X,\mathcal{A}\otimes\mathcal{K})$ is a stable coarse $X$-algebra;
  2. (2)
    The $C^{*}$-algebra $C_{0}(X,\mathcal{A}\otimes\mathcal{K})$ consisting of all bounded, $(\mathcal{A}\otimes\mathcal{K})$-valued functions vanishing at infinity is a stable coarse $X$-algebra.
  3. (3)
    For a subspace $Y\subset X$, define $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$ to be the closure of
    $\Theta[X,\mathcal{A}\otimes\mathcal{K}]:=\{f\in\ell^{\infty}(X,\mathcal{A}\otimes\mathcal{K}):\exists r>0\text{ s.t. }\operatorname{supp}(f)\subset N_{r}(Y)\},$
    where $N_{r}(Y):=\{x\in X:d(x,Y)\leq r\}$ is the $r$-neighborhood of $Y$. Then $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$ is a stable coarse $X$-algebra.
Let $M$ be a proper metric space in the sense that every closed bounded subset is compact. We fix a net $X\subset M$ in $M$. Recall that a subset $X$ is called a net if there exists a positive number $C,D>0$ such that any pair of distinct points in $X$ has distance at least $C$, and $N_{D}(X)=M$, where $N_{D}(X)=\{m\in M:d(m,X)\leq D\}$. We recall the definition of bounded geometry.
Definition 2.3.
A (not necessarily discrete) metric space $M$ is said to have bounded geometry if there exists $\delta>0$ such that for every $R>0$, there exists $K(R)\geq 0$ such that the number of points in a ball $B(x,R)$ with pairwise distance greater that $\delta$ does not exceed $K(R)$.
From the definition of bounded geometry, it is easy to show the following Lemma (c.f. [WY20] Appendix A.1.10).
Lemma 2.4.
If $M$ is a metric space with bounded geometry, then there exist a net $X\subset M$ and a Borel covering $\{B_{x}\}_{x\in X}$ for $M$ such that
  1. (1)
    for each $x$, $B_{x}$ contains a nonempty open subset of $M$,
  2. (2)
    for any $x,y\in X$ with $x\neq y$, we have $B_{x}\cap B_{y}=\emptyset$,
  3. (3)
    for each $x\in X$, we have $x\in B_{x}$,
  4. (4)
    $\sup_{x\in X}{\rm diam}(B_{x})<\infty$.
Let $\mathcal{A}$ be a $C^{*}$-algebra. We first recall the definition of Roe algebras with coefficients in $\mathcal{A}$. Fix a countable dense subset $M_{0}\subset M$ and a separable, infinite-dimensional Hilbert space $\mathcal{H}$. We denote the standard Hilbert $\mathcal{A}$-module by $\mathcal{H}_{\mathcal{A}}:=\mathcal{A}\otimes\mathcal{H}$. We consider a Hilbert $\mathcal{A}$-module $\ell^{2}(M_{0},\mathcal{H}_{\mathcal{A}})$ with the following structure:
$\left\langle\sum_{m\in M_{0}}a_{m},\sum_{m\in M_{0}}b_{x}\right\rangle=\sum_{m\in M_{0}}\left\langle a_{m},b_{m}\right\rangle_{\mathcal{H}_{\mathcal{A}}},$
$\left(\sum_{m\in M_{0}}a_{m}\right)\cdot a=\sum_{m\in M_{0}}(a_{m}a)$
for all $\sum_{m\in M_{0}}a_{m},\sum_{m\in M_{0}}b_{m}\in\ell^{2}(M_{0},\mathcal{H}_{\mathcal{A}})$ and $a\in\mathcal{A}$. Let $B(\ell^{2}(M_{0},\mathcal{H}_{\mathcal{A}}))$ be the algebra of all adjointable $\mathcal{A}$-linear operators on $\ell^{2}(M_{0},\mathcal{H}_{\mathcal{A}})$.
It is obvious that every bounded $\mathcal{A}$-linear operator $T\in B(\ell^{2}(M_{0},\mathcal{H}_{\mathcal{A}}))$ can be expressed as an $(M_{0}\times M_{0})$-matrix $\left(T_{xy}\right)_{x,y\in M_{0}}$, where $T_{xy}\in\mathcal{A}\otimes B(\mathcal{H})$ for all $x,y\in M_{0}$.
Now we are ready to recall some basic concepts in coarse geometry.
Definition 2.5.
Let $T=\left(T_{xy}\right)_{x,y\in M_{0}}$ be an adjointable operator on $\ell^{2}(M_{0},\mathcal{H}_{\mathcal{A}})$.
  1. (1)
    The propagation of $T$ is defined to be
    ${\rm Propagation}(T):=\sup\{d(x,y):T_{x,y}\neq 0\}\in[0,\infty].$
    We say that $T$ has finite propagation if ${\rm Propagation}(T)<\infty$.
  2. (2)
    The operator $T$ is said to be locally compact, if for every bounded subset $B\subset M$, the operator $\left(T_{xy}\right)_{x,y\in B\cap M_{0}}$ is a compact operator on $\ell^{2}(B\cap M_{0},\mathcal{H}_{\mathcal{A}})$.
The concept of Roe algebra was introduced by J. Roe.
Definition 2.6.
Let $M$ be a metric space with bounded geometry, $M_{0}\subset M$ a countable dense subset, and $\mathcal{A}$ a $C^{*}$-algebra. The Roe algebra of $M$ with coefficients in $\mathcal{A}$ is the $C^{*}$-subalgebra $C^{*}(M,\mathcal{A})$ of $B(\ell^{2}(M_{0},\mathcal{A}\otimes\mathcal{H}))$ generated by all locally compact operators with finite propagation. In particular, when $\mathcal{A}=\mathbb{C}$, $C^{*}(M,\mathbb{C})$ is denoted by $C^{*}(M)$.
Next, we recall the definition of twisted Roe algebras. This notion was first introduced by the first-named author together with L. Guo in [DG24], where the coefficient algebras were taken to be coarse algebras. In the present paper, we instead use stable coarse algebras to define twisted Roe algebras. With this modification, the resulting twisted Roe algebras are invariant under coarse equivalences in a certain sense.
We fix a net $X$ of $M$ and a Borel covering $\{B_{x}\}_{x\in X}$ satisfying the four conditions in Lemma 2.4. For every $T=\left(T_{xy}\right)_{x,y\in M_{0}}\in B(\ell^{2}(M_{0},\mathcal{H}_{\mathcal{A}}))$ and every partial translation $v:D\to R$ on $X$, we define a bounded function
$T^{v}:X\to\mathcal{A}\otimes\mathcal{K}$
in the following way. For every $x\in X$, fix a unitary
$U_{x}:\ell^{2}(B_{x}\cap M_{0})\otimes\mathcal{H}\to\mathcal{H}$ (1)
between Hilbert spaces. This induces an isomorphism on the Hilbert $A$-modules $U_{x}\otimes\text{id}_{\mathcal{A}}:\ell^{2}(B_{x}\cap M_{0})\otimes\mathcal{H}_{\mathcal{A}}\to\mathcal{H}_{\mathcal{A}}$. We define
$T^{v}(x):=\left(\text{id}_{\mathcal{A}}\otimes U_{x}\right)\chi_{B_{x}}T\chi_{B_{v(x)}}\left(\text{id}_{\mathcal{A}}\otimes U_{v(x)}\right)\in\mathcal{A}\otimes\mathcal{K}.$
Note that for a stable coarse $X$-algebra $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$ and $T\in B(\ell^{2}(M_{0},\mathcal{H}_{\mathcal{A}}))$, the condition that $T^{v}\in\Gamma(X,\mathcal{A}\otimes\mathcal{K})$ is independent of the choice of unitaries $\{U_{x}\}$ by the third condition in Definition 2.1. For brevity, we omit $\text{id}_{\mathcal{A}}\otimes U_{x}$.
Definition 2.7.
Let $M$ be a metric space with bounded geometry, $X\subset M$ a net in $M$, and $\left\{B_{x}\right\}_{x\in X}$ a Borel cover satisfying the four conditions in Lemma 2.4. Let $\mathcal{A}$ be a $C^{*}$-algebra, and $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$ a coarse $X$-algebra.
  1. (1)
    The algebraic $\Gamma(X,A\otimes\mathcal{K})$-twisted Roe algebra of $M$, denoted by $\mathbb{C}[M;\Gamma(X,\mathcal{A}\otimes\mathcal{K})]$, is the *-subalgebra of the algebraic Roe algebra $\mathbb{C}[M,\mathcal{A}\otimes\mathcal{K}]$ consisting of all operator $T$ satisfying that $T^{v}\in\Gamma(X,\mathcal{A}\otimes\mathcal{K})$ for all partial translation $v:D\to R$ on $X$.
  2. (2)
    The $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$-twisted Roe algebras $C^{*}(M;\Gamma(X,\mathcal{A}\otimes\mathcal{K}))$ is the completion of the *-algebra $\mathbb{C}[M;\Gamma(X,\mathcal{A}\otimes\mathcal{K})]$ under the operator norm on $\ell^{2}(M_{0},\mathcal{H}_{\mathcal{A}})$.
Examples 2.8.
Let $X$ be a metric space with bounded geometry and $\mathcal{A}$ a $C^{*}$-algebra.
  1. (1)
    For the stable coarse $X$-algebra $\ell^{\infty}(X,\mathcal{A}\otimes\mathcal{K})$, we have that
    $C^{*}(X,\ell^{\infty}(X,\mathcal{A}\otimes\mathcal{K}))=C^{*}(X,\mathcal{A}).$
  2. (2)
    For the stable coarse $X$-algebra $C_{0}(X,\mathcal{A}\otimes\mathcal{K})$, we have that
    $C^{*}(X,C_{0}(X,\mathcal{K}))\cong\mathcal{A}\otimes\mathcal{K}.$
  3. (3)
    For a subspace $Y\subset X$ and the stable coarse $X$-algebra $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$ defined in Example 2.2 (3), we have that
    $C^{*}(X;\Gamma(X,\mathcal{A}\otimes\mathcal{K}))\cong C^{*}(Y,\mathcal{A}).$
In [DG24], the first-named author and L. Guo proved that the twisted coarse Baum–Connes conjecture satisfies the coarse imprimitivity. This can be used to compute the $K$-theory of twisted Roe algebras.
We now prove the coarse invariance. In Proposition 2.9, we assume that the coarse equivalence $f:M\to N$ restricts to a bijection between the fixed nets of $X$ and $Y$, and that it preserves the chosen Borel coverings. In this case, for a given a coarse $Y$-algebra $\Gamma(Y,\mathcal{A}\otimes\mathcal{K})$, we consider the coarse $X$-algebra
$f^{*}(\Gamma(Y,\mathcal{A}\otimes\mathcal{K}))=\{\phi\circ f:\varphi\in\Gamma(Y,\mathcal{A}\otimes\mathcal{K})\}$
Then, the proof of coarse invariance is same as that of Roe algebras, and in this case, it is not necessary to require that the coarse algebras are stable. In Theorem 2.14, we prove a more general coarse invariance.
Proposition 2.9.
Let $f:M\to N$ be a coarse equivalence between metric spaces with bounded geometry. Fix a net $X$ in $M$ such that $f$ is one-to-one on $X$ and $Y:=f(X)$ is a net of $Y$, and fix Borel coverings $\{B_{x}\}_{x\in X}$ of $M$ and $\{C_{y}\}_{y\in Y}$ of $N$ satisfying the four conditions in Lemma 2.4 and $f(B_{x})\subset C_{f(x)}$ for all $x\in X$. For any coarse $Y$-algebra $\Gamma(Y,\mathcal{A}\otimes\mathcal{K})$ and the coarse $X$-algebra defined by $\Gamma(X,\mathcal{A}\otimes\mathcal{K}):=f^{*}(\Gamma(Y,\mathcal{A}\otimes\mathcal{K}))$, we have an isomorphism
$C^{*}(M;\Gamma(X,\mathcal{A}\otimes\mathcal{K}))\cong C^{*}(N;\Gamma(Y,\mathcal{A}\otimes\mathcal{K})).$
Proof.
Let $U_{x}:\ell^{2}(B_{x}\cap M_{0},\mathcal{H})\to\mathcal{H}$ and $V_{y}:\ell^{2}(C_{y}\cap N_{0},\mathcal{H})\to\mathcal{H}$ be the fixed unitaries in (1) between Hilbert spaces for all $x\in X$ and $y\in Y$. For each $x\in X$, we denote $W_{x}:=V_{f(x)}^{*}U_{x}:\ell^{2}(B_{x}\cap M_{0},\mathcal{H})\to\ell^{2}(C_{f(x)}\cap N_{0},\mathcal{H})$ and define
$W:=\bigoplus_{x}(W_{x}\otimes\text{id}_{\mathcal{A}}):\ell^{2}(M_{0},\mathcal{H}_{\mathcal{A}})\to\ell^{2}(N_{0},\mathcal{H}_{\mathcal{A}}).$
One can check that for any $T\in B(\ell^{2}(M_{0},\mathcal{H}_{\mathcal{A}}))$ and $x,y\in X$, we have
$\left(V_{f(x)}\otimes\text{id}_{\mathcal{A}}\right)\chi_{C_{f(x)}}\left(WTW^{*}\right)$ $\chi_{C_{f(y)}}\left(V_{f(y)}^{*}\otimes\text{id}_{\mathcal{A}}\right)$
$=\left(V_{f(x)}\otimes\text{id}_{\mathcal{A}}\right)\left(\chi_{B_{x}}T\chi_{B_{y}}\right)\left(V_{f(y)}^{*}\otimes\text{id}_{\mathcal{A}}\right).$
Then it is routine to check $WC^{*}(M;\Gamma(X,\mathcal{A}\otimes\mathcal{K}))W^{*}=C^{*}(N;\Gamma(Y,\mathcal{A}\otimes\mathcal{K}))$. ∎
Next, we show that twisted Roe algebras with stable coarse algebras are independent of the choice of nets and Borel coverings.
We begin by explaining how a coarse equivalence $f:X\to Y$ induces, via pullback, a stable coarse $X$-algebra $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$ from a stable coarse $Y$-algebra $\Gamma(Y,\mathcal{A}\otimes\mathcal{K})$. Then, we use this to formulate the independence of the choice of nets and coarse invariance of twisted Roe algebras. Here it is crucial to assume that coarse algebras are stable (Definition 2.1). Let $X$ and $Y$ be discrete metric spaces with bounded geometry, $f:X\to Y$ a coarse equivalent map, and $\Gamma(Y,\mathcal{A}\otimes\mathcal{K})$ a stable coarse $Y$-algebra. For $\varphi\in\ell^{\infty}(X,\mathcal{A}\otimes\mathcal{K})$ and any subset $\tilde{X}$ of representatives $\{x_{y}:x_{y}\in f^{-1}(y),~y\in f(X)\}$ from each preimage $f^{-1}(y)$, define $\varphi_{\tilde{X}}\in\ell^{\infty}(Y,\mathcal{A}\otimes\mathcal{K})$ by
$\varphi_{\tilde{X}}(y)=\left\{\begin{array}[]{cc}\varphi(x),&y=f(x),x\in\tilde{X},\\ 0,&\text{otherwise.}\end{array}\right$.
Define
[…]
Lemma 2.10.
The set $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$ is a stable coarse $X$-algebra.
Proof.
It is obvious that $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$ is a $C^{*}$-subalgebra of $\ell^{\infty}(X,\mathcal{A}\otimes\mathcal{K})$. For any partial translation $v$ on $X$ and any $\phi\in\Gamma(X,\mathcal{A}\otimes\mathcal{K})$, we show that $v^{*}\phi\in\Gamma(X,\mathcal{A}\otimes\mathcal{K})$. For any subset $\tilde{Y}\subset f(X)$, we fix a set of representatives
$\tilde{X}:=\{x_{y}:x_{y}\in f^{-1}(y),y\in\tilde{Y}\}$
from each preimage $f^{-1}(y)$ for every $y\in\tilde{Y}$. Then for any $y\in\tilde{Y}$ such that $x_{y}\in\operatorname{Dom}(v)$,
$(v^{*}\phi)_{\tilde{X}}(y)=(v^{*}\phi)(x_{y})=\phi(v(x_{y}))$
and $(v^{*}\phi)_{\tilde{X}}(y)=0$ if $y\notin\tilde{Y}$ or $x_{y}\notin\operatorname{Dom}(v)$.
Decompose $\{y\in\tilde{Y}:x_{y}\in\operatorname{Dom}(v)\}=\bigsqcup_{j=1}^{N}A_{j}$ such that for any $j=1,2,\cdots,N$ and for any distinct points $y,y^{\prime}\in A_{j}$, we have that $f(v(x_{y}))\neq f(v(x_{y^{\prime}}))$. Then we can define a partial translations $v_{j}$ on $Y$ such that $\operatorname{Dom}(v_{j})=A_{j}$ and $v_{j}(y)=f(v(x_{y}))$. Define a set
$\tilde{X}_{j}:=\{v(x_{y}):y\in A_{j}\},$
which contains at most one point from each preimage $f^{-1}(y)$. Then for any $y\in A_{j}$,
$v_{j}^{*}(\phi_{\tilde{X}_{j}})(y)=\phi_{\tilde{X}_{j}}(v_{j}(y))=\phi_{\tilde{X}_{j}}(f(v(x_{y})))=\phi(v(x_{y}))$
and $v_{j}^{*}(\phi_{\tilde{X}_{j}})(y)=0$ if $y\notin A_{j}$. Therefore $v^{*}\phi=\sum_{j=1}^{N}v_{j}^{*}(\phi_{\tilde{X}_{j}})\in\Gamma(X,\mathcal{A}\otimes\mathcal{K})$. It is easy to see $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$ is stable if $\Gamma(Y,\mathcal{A}\otimes\mathcal{K})$ is. ∎
By the same technique of decomposition as above, we have the following.
Lemma 2.11.
Let $X$ and $Y$ be metric spaces with bounded geometry, and $\Gamma(Y,\mathcal{A}\otimes\mathcal{K})$ a stable coarse $Y$-algebra. If $f,g:X\to Y$ are coarse equivariant maps which are close in the sense that
$\sup_{x\in X}d(f(x),g(x))<\infty,$
then $f$ and $g$ induce the same coarse $X$-algebra $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$ by the formula (4).
Remark 2.12.
For two coarse equivalences $f:X\to Y$ and $g:Y\to Z$ between discrete metric spaces with bounded geometry. Assume $\Gamma(Z,\mathcal{A}\otimes\mathcal{K})$ is a coarse $Z$-algebra and $\Gamma(Y,\mathcal{A}\otimes\mathcal{K})$ is the coarse $Y$-algebra induced by $g$ via (4). Denote by $\Gamma_{1}(X,\mathcal{A}\otimes\mathcal{K})$ the coarse $X$-algebra induced by $f$ from $\Gamma(Y,\mathcal{A}\otimes\mathcal{K})$ and denote by $\Gamma_{2}(X,\mathcal{A}\otimes\mathcal{K})$ the coarse $X$-algebra induced by $g\circ f$ from $\Gamma(Z,\mathcal{A}\otimes\mathcal{K})$. Then it is easy to see that
$\Gamma_{1}(X,\mathcal{A}\otimes\mathcal{K})=\Gamma_{2}(X,\mathcal{A}\otimes\mathcal{K}).$
Now, we are ready to show that the definition of twisted Roe algebras is independent of the choice of nets and Borel coverings.
Lemma 2.13.
Let $M$ be a metric space with bounded geometry, and let $X$ and $Y$ be nets of $M$ with associated Borel covers $\{B_{x}\}_{x\in X}$ and $\{C_{y}\}_{y\in Y}$ of $M$ satisfying the four conditions in Lemma 2.4. Suppose that $\Gamma(Y,\mathcal{A}\otimes\mathcal{K})$ is a stable coarse $Y$-algebra. If $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$ is a coarse $X$-algebra induced by any coarse equivalence between $X$ and $Y$, which is close to the identity map on $M$, by the formula (4), then we have an equality
$C^{*}(M;\Gamma(X,\mathcal{A}\otimes\mathcal{K}))=C^{*}(M;\Gamma(Y,\mathcal{A}\otimes\mathcal{K}))$
in $B(\ell^{2}(M_{0},\mathcal{H}_{\mathcal{A}}))$.
Proof.
We fix a coarse equivalence $f:X\to Y$, which is close to the identity on $M$. By Remark 2.12, it suffices to show
$\mathbb{C}[M;\Gamma(Y,\mathcal{A}\otimes\mathcal{K})]\subset\mathbb{C}[M;\Gamma(X,\mathcal{A}\otimes\mathcal{K})].$
Let $T\in\mathbb{C}[M;\Gamma(Y,\mathcal{A}\otimes\mathcal{K})]$ and $v$ a partial translation in $X$. We show that the function $\operatorname{Dom}(v)\ni x\mapsto\chi_{B_{x}}T\chi_{B_{v(x)}}\in\mathcal{A}\otimes\mathcal{K}$ is in $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$.
For each $x\in X$, define
$D_{x}:=\bigcup\{C_{y}:C_{y}\cap B_{x}\neq\phi\}\supseteq B_{x}.$
By the third condition in Definition 2.1, it suffices to show that the function
$\operatorname{Dom}(v)\ni x\mapsto\chi_{D_{x}}T\chi_{D_{v(x)}}\in\mathcal{A}\otimes\mathcal{K}$
is in $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$. We label $\{C_{y}:C_{y}\cap B_{x}\neq\phi\}=\{C_{y_{1}^{x}},C_{y_{2}^{x}},\cdots,C_{y_{n_{x}}^{x}}\}$, where $n_{x}$ is uniformly bounded by some $n$. We regard the operator $\chi_{D_{x}}T\chi_{D_{v(x)}}$ is an $n_{x}$-by-$n_{v(x)}$ matrix with entry in $\mathcal{A}\otimes\mathcal{K}$. By the fourth condition of Definition 2.1, it suffices to show that the function $\varphi:\operatorname{Dom}(v)\ni x\mapsto\chi_{C_{y_{i}^{x}}}T\chi_{C_{y_{j}^{v(x)}}}\in\mathcal{A}\otimes\mathcal{K}$ is in $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$ for any $i$ and $j$. For any subset $\tilde{Y}\subset f(X)$, we fix a set
$\tilde{X}=\{x_{y}:x_{y}\in f^{-1}(y),y\in\tilde{Y}\subset f(X)\}$
of representatives of preimages and show that $\varphi_{\tilde{X}}\in\Gamma(Y,\mathcal{A}\otimes\mathcal{K})$. We may assume $\tilde{X}\subset\operatorname{Dom}(v)$. We decompose $\tilde{Y}=\sqcup_{k=1}^{N}A_{k}$ such that:
  • for any distinct $y_{1},y_{2}\in A_{k}$, we have $y_{i}^{x_{y_{1}}}\neq y_{i}^{x_{y_{2}}}$, and
  • for any distinct $y_{1},y_{2}\in A_{k}$, we have $y_{j}^{v(x_{y_{1}})}\neq y_{j}^{v(x_{y_{2}})}$.
Then we can define partial translations $u_{k}$ and $w_{k}$ such that
  • $\operatorname{Dom}(u_{k})=\operatorname{Dom}(w_{k})=A_{k}$, and
  • $u_{k}(y)=y_{i}^{x_{y}}$ and $w_{k}(y)=y_{j}^{v(x_{y})}$.
Then for $y\in A_{k}$, we have that
$\varphi_{\tilde{X}}(y)=\varphi(x_{y})=\chi_{C_{{y_{i}}^{x_{y}}}}T\chi_{C_{y_{j}}^{v(x_{y})}}=\chi_{C_{u_{k}(y)}}T\chi_{C_{w_{k}(y)}}.$
Therefore, $\varphi_{\tilde{X}}=\sum_{k}u_{k}^{*}\left(T^{w_{k}\circ u_{k}^{-1}}\right)\in\Gamma(Y,\mathcal{A}\otimes\mathcal{K})$. ∎
Combining Proposition 2.9 and Lemma 2.13, we can prove the following coarse invariance of twisted Roe algebras.
Theorem 2.14.
Let $M$ and $N$ be metric spaces with bounded geometry, and let $X\subset M$ and $Y\subset N$ be nets of $X$ and $Y$, respectively. Let $f:M\to N$ be a coarse equivalence, and fix Borel covers $\{B_{x}\}_{x\in X}$ and $\{C_{y}\}_{y\in Y}$ satisfying four conditions in Lemma 2.4, together with a coarse equivalence $g:X\to Y$ that is close to $f|_{X}$.
For a given coarse $Y$-algebra $\Gamma(Y,\mathcal{A}\otimes\mathcal{K})$, let $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$ be the coarse $X$-algebra defined via $g$ as in (9). Then there is an isomorphism
$C^{*}(M;\Gamma(X,\mathcal{A}\otimes\mathcal{K}))\;\cong\;C^{*}(N;\Gamma(Y,\mathcal{A}\otimes\mathcal{K})).$
Proof.
We can take nets $Z\subset M$ and $W\subset N$ such that $f:Z\to W$ is bijective. We can take Borel covers $\{B^{\prime}_{z}\}_{z\in Z}$ and $\{C^{\prime}_{w}\}_{w\in W}$ such that $f(B^{\prime}_{z})\subset C^{\prime}_{f(z)}$. Let $\Gamma(W,\mathcal{A}\otimes\mathcal{K})$ be the stable coarse algebra induced by any coarse equivalence $Y\to W$ that is close to the inclusion $Y\hookrightarrow N$. By defining $\Gamma(Z,\mathcal{A}\otimes\mathcal{K}):=f^{*}\Gamma(W,\mathcal{A}\otimes\mathcal{K})$, by Proposition 2.9 we have
$C^{*}(M;\Gamma(Z,\mathcal{A}\otimes\mathcal{K}))\cong C^{*}(N;\Gamma(Y,\mathcal{A}\otimes\mathcal{K})).$
But by Lemma 2.13 and Remark 2.12, we have
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad C^{*}(M;\Gamma(X,\mathcal{A}\otimes\mathcal{K}))=C^{*}(M;\Gamma(Z,\mathcal{A}\otimes\mathcal{K})).\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\qed$
We will use the following theorem to prove the permanence of the twisted coarse Baum–Connes conjecture with respect to stable coarse algebras under passing to subspaces.
Theorem 2.15.
Let $X$ be a metric space with bounded geometry and $Y\subset X$ a subsapce. For a given coarse $Y$-algebra $\Gamma(Y,\mathcal{A}\otimes\mathcal{K})$, there exists a coarse $X$-algebra $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$ such that
$C^{*}(X;\Gamma(X,\mathcal{A}\otimes\mathcal{K}))\cong C^{*}(Y;\Gamma(Y,\mathcal{A}\otimes\mathcal{K})).$
Proof.
For $r>0$, define $N_{r}(Y):=\{x\in X:d(x,Y)\leq r\}$. Fix a coarse equivalence $q_{r}:N_{r}(Y)\to Y$, which fixes every point in $Y$. Then, this map induces a coarse $N_{r}(Y)$-algebra $\Gamma(N_{r}(Y),\mathcal{A}\otimes\mathcal{K})$ by (4), which is independent of the choice of $q_{r}$ by Lemma 2.11. Therefore for $s<r$, we have $\Gamma(N_{s}(Y),\mathcal{A}\otimes\mathcal{K})\subset\Gamma(N_{r}(Y),\mathcal{A}\otimes\mathcal{K})$ and the inductive limit
[…]
is an well-defined coarse $X$-algebra. We claim
[…]
$E_{r_{0}}\subset\bigsqcup_{j=1}^{N}(E(v_{j})\sqcup E(v_{j}^{-1}))$
for some partial translations $v_{1},\cdots,v_{N}$ such that $\operatorname{Dom}(v_{j})\cap\operatorname{Ran}(v_{j})=\emptyset$ for all $j=1,\cdots,N$, where
$E(v_{j}):=\{(x,v_{j}(x))\in X\times X:x\in\operatorname{Dom}(v_{j})\}.$
For $r>0$ such that $T^{v_{j}}\in\Gamma(N_{r}(Y),\mathcal{A}\otimes\mathcal{K})$ for $j=1,\cdots,N$ (such $r$ exists by perturbing by taking arbitrary small perturbation of $T$, if necessary), we have $\operatorname{supp}(T)\subset N_{r}(Y)\times N_{r}(Y)$ and this means $T\in C^{*}(N_{r}(Y);\Gamma(N_{r}(Y),\mathcal{A}\otimes\mathcal{K}))$. The other inclusion is obvious. Therefore by Theorem 2.14, we have
[…]
2.2 The twisted coarse Baum–Connes conjecture
In this subsection, we shall introduce the twisted coarse Baum–Connes conjecture for a metric space with respect to stable coarse algebras.
To introduce the concept of twisted coarse $K$-homology, we shall first recall the definition of Rips complex.
Definition 2.16.
Let $X$ be a metric space with bounded geometry and let $s>0$. The Rips complex at scale $s$, denoted by $P_{s}(X)$, is the simplicial complex whose vertices are $X$, and a subset $\{x_{1},x_{2},\dots,x_{n}\}$ spans a simplex if $d(x_{i},x_{j})\leq s$ for all $1\leq i,j\leq k$.
Each Rips complex $P_{s}(X)$ can be equipped with a spherical metric as follows. The simplex $\left\{\sum_{i=0}^{n}a_{i}x_{i}\right\}$ spanned by $\{x_{0},x_{1}\dots,x_{n}\}$ can be identified with the positive quadrant $\left\{(c_{0},c_{1},\dots,c_{n})\right\}$ equipped with the standard Riemannian metric via the map
$\sum_{i=1}^{n}a_{i}x_{i}\mapsto\left(\frac{a_{0}}{\sum_{i=0}^{n}a_{i}^{2}},\frac{a_{1}}{\sum_{i=0}^{n}a_{i}^{2}},\dots,\frac{a_{n}}{\sum_{i=0}^{n}a_{i}^{2}}\right).$
Next, the Rips complex $P_{s}(X)$ is equipped with the path metric, and the distance between different components is defined to be infinity. For each $s>0$, we consider a countable dense subset $P_{r}(X)_{0}\subset P_{r}(X)$ such that $P_{s}(X)_{0}\subset P_{s^{\prime}}(X)_{0}$ if $s<s^{\prime}$.
Let $M$ be a metric space with bounded geometry, $X\subset M$ a net and $\{B_{x}\}_{x\in X}$ a Borel cover of $M$ which satisfies the four conditions in Lemma 2.4. For a $C^{*}$-algebra $\mathcal{A}$, we consider a stable coarse $X$-algebra $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$. Now we are ready to define the twisted localization algebras.
Definition 2.17.
The $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$-twisted algebraic localization algebra, which is denoted by $\mathbb{C}_{L}[M;\Gamma(X,\mathcal{A}\otimes\mathcal{K})]$, is defined to be the *-algebra of all uniformly bounded and uniformly continuous maps $f:[1,\infty)\to\mathbb{C}[M;\Gamma(X,\mathcal{A}\otimes\mathcal{K})]$ satisfying
${\rm Propagation}(f(t))\to 0,~\mbox{as}~t\to\infty.$
The $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$-twisted localization algebra $C^{*}_{L}(M,\Gamma(X,\mathcal{A}\otimes\mathcal{K}))$ is the completion of the *-algebra $\mathbb{C}_{L}[M;\Gamma(X,\mathcal{A}\otimes\mathcal{K})]$ under the norm
$\|f\|=\sup_{t\in[1,\infty)}\|f(t)\|$
for each $f\in\mathbb{C}_{L}[P_{r}(X);\Gamma(X,\mathcal{A}\otimes\mathcal{K})]$.
For each $s>0$, there is a natural evaluation-at-one map
$e:C^{*}_{L}(P_{s}(X),\Gamma(X,\mathcal{A}\otimes\mathcal{K}))$ $\to C^{*}(P_{s}(X),\Gamma(X,\mathcal{A}\otimes\mathcal{K}))$
$f$ $\mapsto f(1).$
This is a *-homomorphism and it induces a homomorphism
$e_{*}:K_{*}(C^{*}_{L}(P_{s}(X),\Gamma(X,\mathcal{A}\otimes\mathcal{K})))\to K_{*}(C^{*}(P_{s}(X),\Gamma(X,\mathcal{A}\otimes\mathcal{K}))).$ (6)
Since the right hand side of (6) is independent of $s$, passing to the limit yields the twisted coarse assembly map
$\mu:\lim_{s\to\infty}K_{*}(C^{*}_{L}(P_{s}(X),\Gamma(X,\mathcal{A}\otimes\mathcal{K})))\to K_{*}(C^{*}(X,\Gamma(X,\mathcal{A}\otimes\mathcal{K}))).$
Definition 2.18 (The twisted coarse Baum-Connes conjecture with respect to any stable coarse algebras).
A discrete metric space $X$ with bounded geometry is said to satisfy the twisted coarse conjecture with respect to any stable coarse algebras if the twisted coarse Baum–Connes assembly map
$\mu:\lim_{s\to\infty}K_{*}(C^{*}_{L}(P_{s}(X),\Gamma(X,\mathcal{A}\otimes\mathcal{K})))\to K_{*}(C^{*}(X,\Gamma(X,\mathcal{A}\otimes\mathcal{K}))),$
is an isomorphism for any stable coarse algebra $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$. The kernel of the evaluation map $e$ is denoted by $C^{*}_{L,0}(P_{r}(X),\Gamma(X,\mathcal{A}\otimes\mathcal{K}))$ and is called the obstruction algebra. Then by the six-term exact sequence of $K$-theory of $C^{*}$-algebras, $\mu$ is isomorphic if and only if
$\lim_{s\to\infty}K_{*}(C^{*}_{L,0}(P_{s}(X),\Gamma(X,\mathcal{A}\otimes\mathcal{K})))=0.$
For a metric space with bonded geometry, if it admits a coarse embedding into Hilbert space, then it satisfies the twisted coarse Baum–Connes conjecture with respect to any stable coarse algebra by [DG24, Theorem 3.7]. In this paper, we shall study the twisted coarse Baum–Connes conjecture with respect to any coarse algebras for relative hyperbolic groups.
2.3 Permanence properties of the twisted coarse Baum–Connes conjecture with respect to stable coarse algebras
In this subsection we prove the permanence properties of the twisted coarse Baum–Connes conjecture with respect to any stable coarse algebras under coarse invariance and taking subspaces.
First, we discuss how a continuous and coarsely equivalent map $f:M\to N$ induces a map on the $K$-theory of their twisted localization algebras and show the coarse invariance of the twisted coarse Baum–Connes conjecture. We need the following terminology.
Definition 2.19.
For a map $f:M\to N$ and $\varepsilon>0$, an isometry $V:\ell^{2}(M_{0};\mathcal{H})\to\ell^{2}(N_{0};\mathcal{H})$ between Hilbert spaces is called an $\varepsilon$-cover of $f$ if
$\operatorname{supp}(V)\subset\{(x,y)\in M\times N:~d(f(x),y)\leq\varepsilon\}.$
For a $C^{*}$-algebra $\mathcal{A}$, an $\varepsilon$-cover of $f$ is an isometry $V\otimes\text{id}_{\mathcal{A}}:\ell^{2}(M_{0};\mathcal{H}_{\mathcal{A}})\to\ell^{2}(N_{0};\mathcal{H}_{\mathcal{A}})$ for some $\varepsilon$-cover $V:\ell^{2}(M_{0};\mathcal{H})\to\ell^{2}(N_{0};\mathcal{H})$ between Hilbert spaces.
A uniformly continuous family $\{V_{t}\}_{t\in[1,\infty)}$ of isometries from $\ell^{2}(M_{0};\mathcal{H}_{\mathcal{A}})$ to $\ell^{2}(N_{0};\mathcal{H}_{\mathcal{A}})$ is called an continuous cover of $f$ if there exists a decreasing sequence $\{\varepsilon_{k}\}_{k}$ of positive numbers with $\lim_{k}\varepsilon_{k}=0$ such that $V_{t}$ is an $\varepsilon_{k}$ cover of $f$ for all $t\geq k$.
Let $M$ and $N$ be metric spaces with bounded geometry. Fix their net $X\subset M$ and $Y\subset N$ and Borel covers $\{B_{x}\}_{x\in X}$ and $\{C_{y}\}_{y\in Y}$ of $M$ and $N$ satisfying the four conditions in Lemma 2.4, respectively. Let $f:M\to N$ be a continuous coarsely equivariant map. Fix any stable coarse $Y$-algebra $\Gamma(Y,\mathcal{A}\otimes\mathcal{K})$ and $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$ is the coarse $X$-algebra induced by $f$. Note that we can also take nets $X^{\prime}\subset M$ and $Y^{\prime}\subset N$ and Borel covers $\{B^{\prime}_{x}\}_{x\in X^{\prime}}$ and $\{C^{\prime}_{y}\}_{y\in Y^{\prime}}$ of $M$ and $N$, respectively such that $f$ is a bijection between $X^{\prime}$ and $Y^{\prime}$ and $f(B^{\prime}_{x})\subset C^{\prime}_{f(x)}$. We fix a decreasing sequence $\{\varepsilon_{j}\}_{j}$ of positive numbers with $\lim_{j}\varepsilon_{j}=0$. For each $x\in X^{\prime}$, since $f$ is continuous, by decomposing $C^{\prime}_{f(x)}$ into a smaller disjoint Borel sets satisfying the condition in Lemma 2.4, we obtain a unitary
$V_{j}(x):\ell^{2}(B^{\prime}_{x}\cap M_{0},\mathcal{H})\to\ell^{2}(C^{\prime}_{f(x)}\cap N_{0},\mathcal{H})$
such that $\operatorname{supp}V_{j}(x)\subset\{(p,q)\in B^{\prime}_{x}\times C^{\prime}_{f(x)}:d(f(p),q)\leq\varepsilon_{j}\}$. We denote
$V_{j}:=\bigoplus_{x\in X^{\prime}}V_{j}(x)\otimes\text{id}_{\mathcal{A}}:\ell^{2}(M_{0},\mathcal{H}_{\mathcal{A}})\to\ell^{2}(N_{0},\mathcal{H}_{\mathcal{A}}),$
then $V_{j}$ is an $\varepsilon_{j}$-cover of $f$. We denote by $\Gamma(X^{\prime},\mathcal{A}\otimes\mathcal{K})$ and $\Gamma(Y^{\prime},\mathcal{A}\otimes\mathcal{K})$ the coarse algebra induced by the coarse equivalence between $X$ and $X^{\prime}$, and between $Y$ and $Y^{\prime}$, respectively. Then it is easy to see that the adjoint by $V_{j}$ sends $C^{*}(M;\Gamma(X^{\prime},\mathcal{A}\otimes\mathcal{K}))$ to $C^{*}(N;\Gamma(Y^{\prime},\mathcal{A}\otimes\mathcal{K}))$. But by Lemma 2.13, for originally fixed $\{B_{x}\}_{x\in X}$ and $\{C_{y}\}_{y\in Y}$, the adjoint by $V_{j}$ sends $C^{*}(M;\Gamma(X,\mathcal{A}\otimes\mathcal{K}))$ to $C^{*}(N;\Gamma(Y,\mathcal{A}\otimes\mathcal{K}))$. Define a uniformly continuous family of unitaries $(W(t))_{t\in[1,\infty)}$, where
$W(t):\ell^{2}\left(M_{0},\mathcal{H}\otimes\mathbb{C}^{2}\otimes\mathcal{A}\right)\to\ell^{2}\left(N_{0},\mathcal{H}\otimes\mathbb{C}^{2}\otimes\mathcal{A}\right)$
by
$W(t)=R(t-j+1)(V_{j}\oplus V_{j+1})R(t-j+1)^{*}$ (7)
for all $j-1\leq t\leq j$, where $R(t)$ is the rotation
$R(t)=\begin{pmatrix}\cos{(\pi t/2)}&\sin{(\pi t/2)}\\ -\sin{(\pi t/2)}&\cos{(\pi t/2)}\\ \end{pmatrix}$.
Then $(W(t))_{t}$ is a continuous cover of $f$ and the adjoint by $(W(t))_{t}$ intertwines $C^{*}_{L}(M;\Gamma(X,\mathcal{A}\otimes\mathcal{K}))$ and $C^{*}_{L}(N;\Gamma(Y,\mathcal{A}\otimes\mathcal{K}))$. We denote
$f_{*}:=\text{ad}(W(t))_{*}:K_{*}(C^{*}_{L}(M;\Gamma(X,\mathcal{A}\otimes\mathcal{K})))\to K_{*}(C^{*}_{L}(N;\Gamma(Y,\mathcal{A}\otimes\mathcal{K}))).$
Lemma 2.20.
The map $f_{*}$ is independent of the choice of cover $\{W_{t}\}$ of $f$ such that
$\text{ad}(W(t))\left(C^{*}_{L}(M;\Gamma(X,\mathcal{A}\otimes\mathcal{K}))\right)\subset C^{*}_{L}(N;\Gamma(Y,\mathcal{A}\otimes\mathcal{K})).$
Proof.
Assume that families of isometries $\{W_{1}(t)\}$ and $\{W_{2}(t)\}$ are covers of $f$. Since the unitary
$U(t):=\begin{pmatrix}1-W_{1}(t)W_{1}(t)^{*}&W_{1}(t)W_{2}(t)^{*}\\ W_{2}(t)W_{1}(t)^{*}&1-W_{2}(t)W_{2}(t)^{*}\\ \end{pmatrix}$
intertwines $\begin{pmatrix}W_{1}(t)TW_{1}(t)^{*}&0\\ 0&0\\ \end{pmatrix}$ and $\begin{pmatrix}0&0\\ 0&W_{2}(t)TW_{2}(t)^{*}\\ \end{pmatrix}$, it suffices to show that the function $\{U(t)\}$ is in the multiplier of $C^{*}_{L}(N;\Gamma(Y,\mathcal{A}\otimes\mathcal{K}))\otimes M_{2}$. It is easy to see that $\{W_{i}(t)W_{j}(t)^{*}\}$ is in the multiplier of $C^{*}_{L}(N)$ for $i,j=1,2$. For every partial translation $v$ on $Y$, $(T_{t})\in C^{*}_{L}(N;\Gamma(Y,\mathcal{A}\otimes\mathcal{K}))$ and $t\in[1,\infty)$, it suffices to show that the map $\left(T_{t}W_{i}(t)W_{j}(t)^{*}\right)^{v}:Y\to\mathcal{A}\otimes\mathcal{K}$ is in $\Gamma(Y,\mathcal{A}\otimes\mathcal{K})$. We have that
$\left(T_{t}W_{i}(t)W_{j}(t)^{*}\right)^{v}(y)$ $=\chi_{C_{y}}\left(T_{t}W_{i}(t)W_{j}(t)^{*}\right)\chi_{C_{v}(y)}$
$=\sum_{z}\left(\chi_{C_{y}}T_{t}\chi_{C_{z}}\right)\left(\chi_{C_{z}}W_{i}(t)W_{j}(t)^{*}\chi_{C_{v}(y)}\right)$
$=\sum_{z}\left(\chi_{C_{y}}T_{t}\chi_{C_{z}}\right)\left(\chi_{C_{z}}(W^{\prime}_{i}(t)W^{\prime}_{j}(t)^{*}\otimes\text{id}_{\mathcal{A}})\chi_{C_{v(y)}}\right)$
for some covers $\{W_{i}^{\prime}(t)\}$ and $\{W_{j}^{\prime}(t)\}$ between Hilbert spaces $\ell^{2}(M_{0},\mathcal{H})$ and $\ell^{2}(N_{0},\mathcal{H})$. By the third condition in Definition 2.1, it is easy to see that $\left(T_{t}W_{i}(t)W_{j}(t)^{*}\right)^{v}$ is in $\Gamma(Y,\mathcal{A}\otimes\mathcal{K})$. ∎
Next, we discuss the coarse invariance of the twisted coarse Baum–Connes conjecture with respect to stable coarse algebras. The following definition is from [Yu97].
Definition 2.21.
Let $f,g:M\to N$ be continuous maps between metric spaces $M$ and $N$ with bounded geometry. A continuous homotopy $\{F(t,\cdot)\}_{t\in[0,1]}$ between $f$ and $g$ is said to be strongly Lipschitz if
  1. (1)
    for each $t$, $F(t,\cdot):X\to Y$ is a continuous map;
  2. (2)
    there exists a constant $C>0$ such that $d(F(t,x),F(t,y))\leq Cd(x,y)$ for all $x,y\in X$ and $t\in[0,1]$;
  3. (3)
    the family $\{F(\cdot,x)\}_{x\in X}$ is uniformly equicontinuous;
  4. (4)
    $F(0,\cdot)=f$, $F(1,\cdot)=g$.
Lemma 2.22.
Let $X$ and $Y$ be metric spaces with bounded geometry such that there exists a coarse equivalence $f:X\to Y$ and $g:Y\to X$. For any coarse $Y$-algebra $\Gamma(Y,\mathcal{A}\otimes\mathcal{K})$, if $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$ is the coarse $X$-algebra defined by the formula (4) along the coarse equivalence $f$, then we have that
[…]
Proof.
[…]
$V_{f}(t)$ $:\ell^{2}(P_{s}(X)_{0},\mathcal{H}_{\mathcal{A}})\to\ell^{2}(P_{s^{\prime}}(Y)_{0},\mathcal{H}_{\mathcal{A}})$
$V_{g}(t)$ $:\ell^{2}(P_{s^{\prime}}(Y)_{0},\mathcal{H}_{\mathcal{A}})\to\ell^{2}(P_{s^{\prime\prime}}(X)_{0},\mathcal{H}_{\mathcal{A}})$
which intertwines the twisted localization algebras. Since the family $\{V_{g}(t)V_{f}(t)\}_{t}$ covers $g\circ f$, it suffices to show that
$(g\circ f)_{*}={\iota_{s,s^{\prime\prime}}}_{*}:K_{*}(C^{*}_{L}(P_{s}(X);\Gamma(X,\mathcal{A}\otimes\mathcal{K})))\to K_{*}(C^{*}_{L}(P_{s^{\prime\prime}}(X);\Gamma(X,\mathcal{A}\otimes\mathcal{K}))),$
where $\iota_{s,s^{\prime\prime}}$ is the natural inclusion $P_{s}(X)\to P_{s^{\prime\prime}}(X)$. Let $F(t,x)$ be a strong Lipschitz homotopy between $g\circ f$ and $\iota_{s,s^{\prime\prime}}$ ($F(0,\cdot)=g\circ f$ and $F(1,\cdot)=\iota_{s,s^{\prime\prime}}$). Take a double sequence $\{t_{i,j}\}_{i,j}$ of real numbers in $[0,1]$, a decreasing sequence $\{\varepsilon_{j}\}$ of positive numbers such that $\lim\varepsilon_{j}=0$ and a sequence of natural numbers $\{N_{j}\}$ such that
  1. (1)
    $\lim_{j}\varepsilon_{j}=0$, $t_{0,j}=0$ for all $j$ and $t_{i,j}=1$ if $i\geq N_{j}$;
  2. (2)
    $d(F(t_{i,j},\cdot)(x),F(t_{i+1,j},\cdot)(x))\leq\varepsilon_{j}$;
  3. (3)
    $d(F(t_{i,j},\cdot)(x),F(t_{i,j+1},\cdot)(x))\leq\varepsilon_{j}$.
For each $i$ and $j$, take $\varepsilon_{j}$-covers $V_{i,j}$ of $F(t_{i,j},\cdot)$ and define a family $\{W_{i}(t)\}_{t\in[1,\infty)}$ by the same formula as (7). It suffices to show that $\text{ad}(W_{0})(u)]=[u]\in K_{1}(C^{*}_{L}(P_{s^{\prime\prime}}(X);\Gamma(X,\mathcal{A}\otimes\mathcal{K})))$ for any unitary $u$ in the unitaization of $C^{*}_{L}(P_{s}(X);\Gamma(X,\mathcal{A}\otimes\mathcal{K}))$. (We extend $u$ to the identity on the complement $\ell^{2}(P_{s^{\prime\prime}}(X)_{0}\setminus P_{s}(X)_{0},\mathcal{H}_{\mathcal{A}})$ to regard it as an unitary in the unitaization of $C^{*}_{L}(P_{s^{\prime\prime}}(X);\Gamma(X,\mathcal{A}\otimes\mathcal{K}))$. Define unitaries $a,b,c$ acting on $\ell^{2}(P_{s^{\prime\prime}}(X)_{0},\ell^{2}\otimes\mathcal{H}_{\mathcal{A}})$ by
$a$ $=\bigoplus_{i=0}^{\infty}\text{ad}(W_{i})(u)u^{*}$
$b$ $=\bigoplus_{i=0}^{\infty}\text{ad}(W_{i+1})(u)u^{*}$
$c$ $=I\oplus\bigoplus_{i=1}^{\infty}\text{ad}(W_{i})(u)u^{*}.$
[…]
Combining this with Theorem 2.14, we obtain the following permanence result.
Corollary 2.23.
Let $X$ and $Y$ be metric space with bounded geometry which are coarsely equivalent to each other. Then $X$ satisfies the twisted coarse Baum–Connes conjecture with respect to any coarse algebras if and only if $Y$ does.
Next, we prove the permanence of the twisted coarse Baum–Connes conjecture with respect to any stable coarse algeras under taking subspaces.
Theorem 2.24.
Let $Y\subset X$ be metric spaces with bounded geometry. For any stable coarse $Y$-algebra $\Gamma(Y,\mathcal{A}\otimes\mathcal{K})$, let $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$ be the stable coarse $X$-algebra defined in (5). Then the natural inclusion
$C^{*}_{L}(P_{s}(Y);\Gamma(Y,\mathcal{A}\otimes\mathcal{K}))\to C^{*}_{L}(P_{s}(X);\Gamma(X,\mathcal{A}\otimes\mathcal{K}))$
induces an isomorphism
[…]
Moreover, if $X$ satisfies the twisted coarse Baum–Connes conjecture with respect to any stable coarse algebras then $Y$ does.
Before proving this, we need following preliminary Definition and Lemma.
Definition 2.25.
Let $M$ be any metric space with bounded geometry and a net $X$ and a subset $Y\subset X$. Let $\Gamma(Y,\mathcal{A}\otimes\mathcal{K})$ be any stable coarse $Y$-algebra and $\Gamma(N_{r}(Y),\mathcal{A}\otimes\mathcal{K})$ and $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$ are defined in the proof of Theorem 2.15. We define a $C^{*}$-algebra $C^{*}_{L,\text{uniform}}(M;\Gamma(X,\mathcal{A}\otimes\mathcal{K}))$ to be the closure of the *-algebra consisting all $(T_{t})_{t}$ in $\mathbb{C}_{L}[M;\Gamma(X,\mathcal{A}\otimes\mathcal{K})]$ such that for any partial translation $v$ there exists $r>0$ such that $(T_{t})^{v}\in\Gamma(N_{r}(Y),\mathcal{A}\otimes\mathcal{K})$ for all $t\in[1,\infty)$.
Lemma 2.26.
Under the setting of Definition 2.25, if $M$ is a finite-dimensional simplicial complex, then we have an isomorphism
$K_{*}(C^{*}_{L,\text{uniform}}(M;\Gamma(X,\mathcal{A}\otimes\mathcal{K})))\cong K_{*}(C^{*}_{L}(M;\Gamma(X,\mathcal{A}\otimes\mathcal{K})))$
induced by the natural inclusion $C^{*}_{L,\text{uniform}}(M;\Gamma(X,\mathcal{A}\otimes\mathcal{K}))\to C^{*}_{L}(M;\Gamma(X,\mathcal{A}\otimes\mathcal{K}))$.
Proof.
We prove this by the induction on the dimension $m$ of $M$. When $m=0$, in this case we may assume $M=X$ and it suffices to show that the inclusion
[…]
induces an isomorphism on $K$-theory, where for any $C^{*}$-algebra $\mathcal{B}$, $C_{ub}([1,\infty),\mathcal{B})$ is the $C^{*}$-algebra of all uniformly continuous bounded function from $1,\infty)$ to $\mathcal{B}$. But since $\Gamma(N_{r}(Y),\mathcal{A}\otimes\mathcal{K}))$ and $\Gamma(X,\mathcal{A}\otimes\mathcal{K}))$ are quasi-stable, the evaluation maps at 11
$C_{ub}([1,\infty),\Gamma(N_{r}(Y),\mathcal{A}\otimes\mathcal{K}))$ $\to\Gamma(N_{r}(Y),\mathcal{A}\otimes\mathcal{K})$
$C_{ub}([1,\infty),\Gamma(X,\mathcal{A}\otimes\mathcal{K}))$ $\to\Gamma(X,\mathcal{A}\otimes\mathcal{K}))$
induces isomorphisms on $K$-theory by [WY20, Lemma 12.4.3]. The assertion follows from the continuity of $K$-theory. Next, we assume induction hypothesis for the dimension $0,\cdots,m-1$. For each $m$-dimensional simplicial complex $\Delta$, we denote by $c(\Delta)$ the center of it. We define
$\Delta_{1}:=\left\{x\in\Delta:d(x,c(\Delta))\leq\frac{1}{10}\right\},\quad\Delta_{2}:=\left\{x\in\Delta:d(x,c(\Delta))\geq\frac{1}{10}\right\}$
and define
$N_{1}^{\prime}:=\bigcup\{\Delta_{1}$ $:\Delta\text{ is an $m$-dimensional simplex of }M\}$
[…]
$N_{2}^{\prime}:=\bigcup\{\Delta_{2}$ $:\Delta\text{ is an $m$-dimensional simplex of }M\}$
[…]
By taking a point from each simplex, we fix a net $X^{\prime}_{1}\subset N_{1}^{\prime}$, $X^{\prime}_{2}\subset N_{2}^{\prime}$ and $X_{1,2}^{\prime}\subset N^{\prime}_{1}\cap N^{\prime}_{2}$, respectively. Since $N_{1}^{\prime}$, $N_{2}^{\prime}$ and $N_{1}^{\prime}\cap N_{2}^{\prime}$ are strongly Lipschitz homotopic to simplicial complexes $N_{1}$, $N_{2}$ and $N_{1,2}$ whose dimensions are less than $m$, respectively through coarse equivalent maps. We denote the corresponding nets $X_{1}\subset N_{1}$, $X_{2}\subset N_{2}$ and $X_{1,2}\subset N_{1,2}$, respectively. By the same argument as Lemma 2.22, one can show that
$K_{*}(C^{*}_{L,\text{uniform}}(N_{i}^{\prime};\Gamma(X_{i}^{\prime},\mathcal{A}\otimes\mathcal{K})))$ $\cong K_{*}(C^{*}_{L,\text{uniform}}(N_{i};\Gamma(X_{i},\mathcal{A}\otimes\mathcal{K})))$
$K_{*}(C^{*}_{L}(N_{i}^{\prime};\Gamma(X_{i}^{\prime},\mathcal{A}\otimes\mathcal{K})))$ $\cong K_{*}(C^{*}_{L}(N_{i};\Gamma(X_{i},\mathcal{A}\otimes\mathcal{K})))$
for $i=1,2$ and
$K_{*}(C^{*}_{L,\text{uniform}}(N_{1}^{\prime}\cap N_{2}^{\prime};\Gamma(X_{1,2}^{\prime},\mathcal{A}\otimes\mathcal{K})))$ $\cong K_{*}(C^{*}_{L,\text{uniform}}(N_{1,2};\Gamma(X_{1,2},\mathcal{A}\otimes\mathcal{K})))$
$K_{*}(C^{*}_{L}(N_{1}^{\prime}\cap N_{2}^{\prime};\Gamma(X_{1,2}^{\prime},\mathcal{A}\otimes\mathcal{K})))$ $\cong K_{*}(C^{*}_{L}(N_{1,2};\Gamma(X_{1,2},\mathcal{A}\otimes\mathcal{K}))).$
Note that $K$-theory for both $C^{*}_{L,\text{uniform}}(M;\Gamma(X,\mathcal{A}\otimes\mathcal{K})))$ and $C^{*}_{L}(M;\Gamma(X,\mathcal{A}\otimes\mathcal{K})))$ admits Mayer-Vietoris sequence by the same proof as [Yu97, Proposition 3.11]. Therefore the assertion follows from the five lemma and inductive hypothesis. ∎
Proof of Theorem 2.24.
We construct a map
[…]
which is the inverse of a map induced by the inclusion
$C^{*}_{L}(P_{s}(Y);\Gamma(Y,\mathcal{A}\otimes\mathcal{K}))\to C^{*}_{L,\text{uniform}}(P_{s}(X);\Gamma(X,\mathcal{A}\otimes\mathcal{K})).$ (8)
Take any unitary in the unitaization of $C^{*}_{L,\text{uniform}}(P_{s}(X);\Gamma(X,\mathcal{A}\otimes\mathcal{K}))$. We may assume that there exists $R\geq 0$ and a scalar $\lambda\in\mathbb{C}$ such that $u_{t}-\lambda$ is supported in $(P_{s}(N_{R}(Y)))\times(P_{s}(N_{R}(Y)))$ for every $t$. Therefore we can regard $u_{t}]$ as an element in $K_{*}(C^{*}_{L}(P_{s}(N_{R}(Y));\Gamma(Y,\mathcal{A}\otimes\mathcal{K})))$. But since we have an isomorphism
[…]
2.4 Family of subspaces
In this subsection, for a given metric space $X$ with bounded geometry and a family of subspaces $\{Y^{(j)}\}$ of $X$, we study the connection between the twisted coarse Baum–Connes conjecture of the family $\{Y^{(j)}\}$ and that of the space $X$.
We first realize the $K$-theory of twisted Roe algebras of $\bigsqcup_{j}Y^{(j)}$ as that of $X$ by changing the coefficient algebra $\mathcal{A}$. We denote by $\tilde{d}$ the distance on $\bigsqcup_{j}Y^{(j)}$ defined by
$\tilde{d}(x,y)=\left\{\begin{array}[]{cc}d(x,y)&x,y\in Y^{(j)}\text{ for some }j\\ \infty&\text{otherwise.}\end{array}\right$.
[…]
$C^{*}(\bigsqcup_{j}X;\Gamma(\bigsqcup_{j}X,\mathcal{A}\otimes\mathcal{K}))\cong C^{*}(\bigsqcup_{j}Y^{(j)};\Gamma(\bigsqcup_{j}Y^{(j)},\mathcal{A}\otimes\mathcal{K})),$
where $\Gamma(\bigsqcup_{j}X,\mathcal{A}\otimes\mathcal{K})$ is constructed from the coarse $\bigsqcup Y^{(j)}$-algebra $\Gamma(\bigsqcup_{j}Y^{(j)};\mathcal{A}\otimes\mathcal{K})$ and the inclusion $\bigsqcup Y^{(j)}\subset\bigsqcup X$ by the formula (5).
Remark 2.27.
Let $\mathcal{A}$ be a $C^{*}$-algebra, $M$ a metric space with bounded geometry and a countable dense subset $M_{0}$ and a net $X\subset M$. For isometries $V:\mathcal{H}_{\mathcal{A}}\to\ell^{2}\otimes\mathcal{H}_{\mathcal{A}}\cong\mathcal{H}_{\mathcal{A}},~\xi\mapsto(\xi,0,0,\cdots)$, and $W:\ell^{2}\otimes\mathcal{H}_{\mathcal{A}}\cong\mathcal{H}_{\mathcal{A}}\to\ell^{2}\otimes\mathcal{H}_{\mathcal{A}},~\xi\mapsto(\xi,0,0,\cdots)$, we denote the isometry induced on the geometric modules
$\tilde{V}:\ell^{2}(M_{0},\mathcal{H}_{\mathcal{A}})$ $\to\ell^{2}(M_{0},\mathcal{H}_{\mathcal{A}})$
$\tilde{W}:\ell^{2}(M_{0},\ell^{2}\otimes\mathcal{H}_{\mathcal{A}})$ $\to\ell^{2}(M_{0},\ell^{2}\otimes\mathcal{H}_{\mathcal{A}}).$
For any stable coarse $X$-algebra $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$, the adjoint by $\tilde{V}$ and $\tilde{W}$ can be restricted to the twisted Roe algebra $C^{*}(M,\Gamma(X,\mathcal{A}\otimes\mathcal{K}))$ by the fourth condition of Definition 2.1 and the induced map on $K$-theory is the identity map by the same arguements in the proof as [WY20, Lemma 5.1.12].
Lemma 2.28.
Given a coarse $\bigsqcup_{j}X$-algebra $\Gamma(\bigsqcup_{j}X,\mathcal{A}\otimes\mathcal{K})$, we define a new stable coarse $X$-algebra with respect to $\ell^{\infty}(\mathbb{N},\mathcal{A}\otimes\mathcal{K})\otimes\mathcal{K}$ as
$\Gamma(X;\ell^{\infty}(\mathbb{N},\mathcal{A}\otimes\mathcal{K})\otimes\mathcal{K}):=\left\{f\in\ell^{\infty}(X,\ell^{\infty}(\mathbb{N},\mathcal{A}\otimes\mathcal{K})\otimes\mathcal{K}):\left(\psi(f_{\bullet}(j))\right)_{j}\in\Gamma\left(\bigsqcup_{j}X,\mathcal{A}\otimes\mathcal{K}\right)\right\},$
where $\psi:\mathcal{A}\otimes\mathcal{K}\otimes\mathcal{K}\cong\mathcal{A}\otimes\mathcal{K}$ is the fixed isomorphism. Then we have an isomorphism for Roe algebras and localization algebras on $K$-theory:
$K_{*}\left(C^{*}\left(\bigsqcup_{j}M;\Gamma\left(\bigsqcup_{j}X,\mathcal{A}\otimes\mathcal{K}\right)\right)\right)$ $\cong K_{*}\left(C^{*}\left(M;\Gamma(X;\ell^{\infty}(\mathbb{N},\mathcal{A}\otimes\mathcal{K})\otimes\mathcal{K})\right)\right)$
$K_{*}\left(C_{L}^{*}\left(\bigsqcup_{j}M;\Gamma\left(\bigsqcup_{j}X,\mathcal{A}\otimes\mathcal{K}\right)\right)\right)$ $\cong K_{*}\left(C_{L}^{*}\left(M;\Gamma(X;\ell^{\infty}(\mathbb{N},\mathcal{A}\otimes\mathcal{K})\otimes\mathcal{K})\right)\right).$
Proof.
Define *-homomorphisms
$\alpha:$ $C^{*}\left(\bigsqcup_{j}X;\Gamma\left(\bigsqcup_{j}X,\mathcal{A}\otimes\mathcal{K}\right)\right)\to C^{*}\left(X;\Gamma(X,\ell^{\infty}(\mathbb{N},\mathcal{A}\otimes\mathcal{K})\otimes\mathcal{K})\right)$
$\beta:$ $C^{*}\left(X;\Gamma(X,\ell^{\infty}(\mathbb{N},\mathcal{A}\otimes\mathcal{K})\otimes\mathcal{K})\right)\to C^{*}\left(\bigsqcup_{j}X;\Gamma\left(\bigsqcup_{j}X,\mathcal{A}\otimes\mathcal{K}\right)\right)$
by
$\alpha((T^{(j)}))_{x,y}$ $:=(j\mapsto T^{(j)}_{x,y})\otimes e_{1,1}$
$\beta(S)^{(j)}_{x,y}$ $:=\psi(S_{x,y}(j)).$
The compositions $\beta\circ\alpha$ and $\alpha\circ\beta$ are the adjoints by $\tilde{V}$ and $\tilde{W}$ in the previous remark, respectively, and both induce isomorphisms on $K$-theory. The same proof works for localization algebras. ∎
Using the above lemma, we are able to reduce the twisted coarse Baum–Connes conjecture with respect to any stable coarse algebras for $\bigsqcup X$ to that of $X$.
3 Filtered $C^{*}$-algebra and controlled $K$-theory
In this section, we recall the concept of filtered $C^{*}$-algebra and controlled $K$-theory from [GWY24]. Using these techniques, we are able to obtain a controlled Mayer–Vietoris argument for the twisted coarse Baum–Connes conjecture.
In order to use $C^{*}$-algebras to encode geometry of metric spaces, we need the concept of filtration.
Definition 3.1.
A filtration on a $C^{*}$-algebra $A$ is a collection of self-adjoint subspaces $(A_{r})_{r\geq 0}$ of $A$ indexed by the non-negative real numbers that satisfies the following properties:
  1. (1)
    if $r_{1}\leq r_{2}$, then $A_{r_{1}}\subset A_{r_{2}}$;
  2. (2)
    for all $r_{1},r_{2}$, we have $A_{r_{1}}\cdot A_{r_{2}}\subset A_{r_{1}\cdot r_{2}}$;
  3. (3)
    the union $\bigcup_{r\geq 0}A_{r}$ is dense in $A$.
In the case of Roe algebras, the filtartion is defined by propagation. We apply this framework to the obstruction algebras $C^{*}_{L,0}(M;\Gamma(X,\mathcal{A}))$ defined in Definition 2.18. For $r>0$, we define
$C^{*}_{L,0}(M;\Gamma(X,\mathcal{A}\otimes\mathcal{K}))_{r}:=\{f\in C^{*}_{L,0}(M;\Gamma(X,\mathcal{A}\otimes\mathcal{K})):\text{propagation}(f(t))<r~\forall t\in[1,\infty)\}.$
Next, we shall recall the quantitative $K$-theory introduced by G. Yu [Yu98] to prove the coarse Baum-Connes conjecture for metric spaces with finite asymptotic dimension.
Definition 3.2.
Let $A$ be a $C^{*}$-algebra. A quasi-projection in $A$ is an element $p\in A$ such that $p=p^{*}$ and $\|p^{2}-p\|<\frac{1}{8}$.
If $S$ is a self-adjoint subspace of $A$, write $\mathrm{M}_{n}(S)$ for the matrices in $\mathrm{M}_{n}(A)$ with all entries coming from $S$, and let $P^{1/8}_{n}(S)$ denote the collection of quasi-projections in $\mathrm{M}_{n}(S)$.
Let $\chi=\chi_{(1/2,\infty)}$ be the characteristic function of the interval $(1/2,\infty)$. Then $\chi$ is continuous on the spectrum of any quasi-projection, and thus there is a well-defined map:
$\kappa:P^{1/8}_{n}(S)\to P_{n}(A),\quad p\mapsto\chi(p),$
where $P_{n}(A)$ denotes the projections in $\mathrm{M}_{n}(A)$.
Definition 3.3.
Let $A$ be a non-unital $C^{*}$-algebra with a filtration $(A_{r})_{r}$. Let $\tilde{A}_{r}$ denote the subspace $A_{r}+\mathbb{C}1$ of $\tilde{A}$.
Using the inclusions
$P^{1/8}_{n}(\tilde{A}_{r})\ni p\mapsto\begin{pmatrix}p&0\\ 0&0\end{pmatrix}\in P^{1/8}_{n+1}(\tilde{A}_{r}),$
we may define the union
$P^{1/8}_{\infty}(\tilde{A}_{r}):=\bigcup_{n=1}^{\infty}P^{1/8}_{n}(\tilde{A}_{r}).$
[…]
Define an equivalence relation on $P^{1/8}_{\infty}(\tilde{A}_{r})\times\mathbb{N}$ by declaring $(p,m)\sim(q,n)$ if there exists a positive integer $k$ and an element $h\in P^{1/8}_{\infty}(C((0,1],\tilde{A}_{r}))$ such that
$h(0)=\begin{pmatrix}p&0\\ 0&1_{n+k}\end{pmatrix},\quad h(1)=\begin{pmatrix}q&0\\ 0&1_{m+k}\end{pmatrix}$.
For $(p,m)\in P^{1/8}_{\infty}(\tilde{A}_{r})\times\mathbb{N}$, denote by $[p,m]$ its equivalence class under $\sim$.
[…]
Finally, define
$K^{r,1/8}_{0}(A):=\left\{[p,m]\in P_{1/8}^{\infty}(\tilde{A}_{r})\times\mathbb{N}\,\middle|\,\mathrm{rank}(\kappa(\rho(p)))=m\right\}/\sim.$
The set $K^{r,1/8}_{0}(A)$ is equipped with an operation defined by
$p,m]+[q,n]:=\left[\begin{pmatrix}p&0\\ 0&q\end{pmatrix},m+n\right].$
Definition 3.4.
Let $A$ be a unital $C^{*}$-algebra. A quasi-unitary in $A$ is an element $u\in A$ such that
$\|1-uu^{*}\|<\tfrac{1}{8}\quad\text{and}\quad\|1-u^{*}u\|<\tfrac{1}{8}.$
If $S$ is a self-adjoint subspace of $A$ containing the unit, write $U^{1/8}_{n}(S)$ for the collection of quasi-unitaries in $M_{n}(S)$.
Note that since $\|1-u^{*}u\|<\tfrac{1}{8}<1$, the element $u^{*}u$ is invertible. Hence, there is a well-defined map
$\kappa:U^{1/8}_{n}(S)\to U_{n}(A),\quad u\mapsto u(u^{*}u)^{-1/2},$
where $U_{n}(A)$ denotes the group of unitaries in $M_{n}(A)$.
Definition 3.5.
Let $A$ be a non-unital $C^{*}$-algebra with a filtration $(A_{r})_{r}$. Using the inclusions
$U^{1/8}_{n}(\tilde{A}_{r})\ni u\mapsto\begin{pmatrix}u&0\\ 0&1\end{pmatrix}\in U^{1/8}_{n+1}(\tilde{A}_{r}),$
we may define the union
$U^{1/8}_{\infty}(\tilde{A}_{r}):=\bigcup_{n=1}^{\infty}U^{1/8}_{n}(\tilde{A}_{r}).$
Define an equivalence relation on $\mathcal{U}_{1/8}^{\infty}(\tilde{A}_{r})$ by declaring $u\sim v$ if there exists an element
$h\in U^{1/8}_{\infty}(C((0,1],\tilde{A}_{r}))\quad\text{such that}\quad h(0)=u\quad\text{and}\quad h(1)=v.$
For $u\in U^{1/8}_{\infty}(\tilde{A}_{r})$, denote by $[u]$ its equivalence class under $\sim$. Finally, define
$K^{r,1/8}_{1}(A):=U^{1/8}_{\infty}(\tilde{A}_{r})\big{/}\sim.$
The set $K^{1/8}_{1}(A)$ is equipped with a binary operation defined by
$u]+[v]:=\left[\begin{pmatrix}u&0\\ 0&v\end{pmatrix}\right$].
The advantage of quantitative $K$-theory is that it is more computable compared with the usual operator $K$-theory. It also has connections with the usual $K$-theory.
Definition 3.6.
Let $A$ be a non-unital $C^{*}$-algebra, and let $S$ be a self-adjoint subspace of $A$. Let $\kappa$ be one of the maps from Definitions 3.2 or Definition 3.4 (it will be clear from context which is meant).
Define maps:
$c_{0}:K^{r,1/8}_{0}(A)\longrightarrow K_{0}(A),\quad[p,m]\mapsto[\kappa(p)]-[1_{m}],$
$c_{1}:K^{r,1/8}_{1}(A)\longrightarrow K_{1}(A),\quad[u]\mapsto[\kappa(u)],$
$c:=c_{0}\oplus c_{1}:K^{r,1/8}_{*}(A)\longrightarrow K_{*}(A).$
We call $c_{0}$, $c_{1}$, and $c$ the comparison maps.
The map $c$ is asymptotically isomorphic in the following sense:
Proposition 3.7 (Proposition 4.9 [GWY24]).
Let $A$ be a non-unital $C^{*}$-algebra with a filtration $(A_{r})_{r}$. Then for any $x\in K_{*}(A)$, there exists $r>0$ and an element $y\in K^{r,\frac{1}{8}}(A)$ such that $C(y)=x$. Moreover if $x,y\in K^{s,\frac{1}{8}}(A)$ for some $s>0$, then there exists $r>0$ such that $x=y$ in $K^{r,\frac{1}{8}}(A)$.
In the following, we shall recall some techniques for the computation of quantitative $K$-theory formulated in [GWY24]. Those techniques are crucial to reduce to computation of the $K$-theory of Roe algebra of relative hyperbolic group to that of peripheral subgroups.
Definition 3.8.
Let $A$ be a filtered $C^{*}$-algebra, and $I$ be a $C^{*}$-ideal in $A$ equipped with its own filtration. We say that $I$ is a filtered ideal of $A$ if for any $r\geq 0$, $I_{r}\subset A_{r}$, and if for any $r,s\geq 0$, $A_{s}\cdot I_{r}\cup I_{r}\cdot A_{s}\subset I_{s+r}$.
Definition 3.9.
Let $(I^{\omega},J^{\omega};A^{\omega})_{\omega\in\Omega}$ be an indexed set of triples, where each $A^{\omega}$ is a filtered $C^{*}$-algebra, and each $I^{\omega}$ and $J^{\omega}$ is a filtered ideal in $A^{\omega}$. Give each stabilization $A^{\omega}\otimes\mathcal{K}$, $I^{\omega}\otimes\mathcal{K}$ and $J^{\omega}\otimes\mathcal{K}$ the filtration
$(A^{\omega}\otimes\mathcal{K})_{r}=A^{\omega}_{r}\otimes\mathcal{K},~(I^{\omega}\otimes\mathcal{K})_{r}=I^{\omega}_{r}\otimes\mathcal{K},~(J^{\omega}\otimes\mathcal{K})_{r}=J^{\omega}_{r}\otimes\mathcal{K}.$
The collection $(I^{\omega},J^{\omega};A^{\omega})_{\omega\in\Omega}$ of pairs of ideals and $C^{*}$-algebras containing them is uniformly excisive if for any $r_{0},m_{0}\geq 0$ and $\epsilon>0$, there are $r\geq r_{0}$, $m\geq 0$, and $\delta>0$ such that:
  • (i)
    for any $\omega\in\Omega$ and any $a\in(A^{\omega}\otimes\mathcal{K})_{r_{0}}$ of norm at most $m_{0}$, there exist elements $b\in(I^{\omega}\otimes\mathcal{K})_{r}$ and $c\in(J^{\omega}\otimes\mathcal{K})_{r}$ of norm at most $m$ such that $\|a-(b+c)\|<\epsilon$;
  • (ii)
    for any $\omega\in\Omega$ and any $a\in I^{\omega}\otimes\mathcal{K}\cap J^{\omega}\otimes\mathcal{K}$ such that
    $d(a,(I^{\omega}\otimes\mathcal{K})_{r_{0}})<\delta\quad\text{and}\quad d(a,(J^{\omega}\otimes\mathcal{K})_{r_{0}})<\delta,$
    there exists $b\in I^{\omega}_{r}\otimes\mathcal{K}\cap J^{\omega}_{r}\otimes\mathcal{K}$ such that $\|a-b\|<\epsilon$.
Theorem 3.10 (Proposition 7.7 [GWY24]).
Let $\{(I^{\omega},J^{\omega};A^{\omega})\}_{\omega\in\Omega}$ be a uniformly excisive collection, where all algebras and ideals are non-unital. Then, for any $r_{0}\geq 0$, there exist $r_{1},r_{2}\geq r_{0}$ satisfying the following:
For each $\omega\in\Omega$ and each $x\in K^{r_{0},1/8}_{*}(A^{\omega})$, there exists an element
$\partial_{c}(x)\in K^{r_{1},1/8}_{*}(I^{\omega}\cap J^{\omega})$
such that if $\partial_{c}(x)=0$, then there exist elements
$y\in K^{r_{2},1/8}_{*}(I^{\omega}),\quad z\in K^{r_{2},1/8}_{*}(J^{\omega})$
such that
$x=y+z\quad\text{in }K^{r_{2},1/8}_{*}(A^{\omega}).$
Furthermore, the boundary map $\partial_{c}$ has the following naturality property:
Let $\{(K^{\theta},L^{\theta};B^{\theta})\}_{\theta\in\Theta}$ be another uniformly excisive collection of non-unital filtered $C^{*}$-algebras and ideals. Assume there exists a map $\pi:\Theta\to\Omega$ and, for each $\theta\in\Theta$, an inclusion
$A^{\pi(\theta)}\subset B^{\theta}$
such that for each $r\geq 0$:
$A^{\pi(\theta)}_{r}\subset B^{\theta}_{r},\quad I^{\pi(\theta)}_{r}\subset K^{\theta}_{r},\quad J^{\pi(\theta)}_{r}\subset L^{\theta}_{r}.$
Let $r_{0}$ be given, and let $r_{1}$ be as in the statement above for both uniformly excisive families. Then the following diagram commutes:
$K^{r_{0},1/8}_{*}(A^{\pi(\theta)})$$K^{r_{1},1/8}_{*}(I^{\pi(\theta)}\cap J^{\pi(\theta)})$$K^{r_{0},1/8}_{*}(B^{\theta})$$K^{r_{1},1/8}_{*}(K^{\theta}\cap L^{\theta}),$$\scriptstyle{\partial_{c}}$$\scriptstyle{\partial_{c}}$
where the vertical maps are induced by subspace inclusions.
4 Relative hyperbolic groups, decomposition complexity and the strategy of the proof
In this section, we recall some properties of relative hyperbolic groups from [Osi05] and outline the strategy of the proof of Theorem 1.3. In this section, we assume that $G$ is a finitely generated group which is hyperbolic relative to a finite family of subgroups $\{H_{1},\cdots,H_{n}\}$. We fix a symmetric generating set $S$ of $G$ and denote by $d_{S}$ and $d_{s\cup\bigcup H_{i}}$ the Cayley distance on $G$ with respect to the generating set $S$ and $S\cup\bigcup H_{i}$, respectively. We denote the ball of radius $n$ with respect to $d_{s\cup\bigcup H_{i}}$ centered at $e$ by $B(n)$.
Lemma 4.1 (Proposition 6.8 of [Osi06]).
In this setting, the metric space $(G,d_{s\cup\bigcup H_{i}})$ is hyperbolic.
The next lemma plays the crucial role in the proof of the twisted Baum–Connes conjecture of $B(n)$. By definition of $B(n)$, we have
$B(n)=\left(\bigcup_{j}B(n-1)H_{j}\right)\cup B(n-1)S.$
Take a representative $R(n-1)\subset B(n-1)$ from a subset $B(n-1)H_{j}\subset G/H_{j}$ of the right cosets so that we have
$B(n-1)H_{j}=\bigsqcup_{g\in R(n-1)}gH_{j}.$
Lemma 4.2 (Proposition 6.1 of [DG07] and proof of Lemma 3.2 of [Osi05]).
For every $L>0$, there exists $\kappa(L)>0$ such that if
$Y=\{x\in G:d_{S}(x,B(n-1))\leq\kappa(L)\}$
then for each $j$,
$B(n-1)H_{j}\subset Y\sqcup\left(\bigsqcup_{g\in R(n-1)}gH_{j}\setminus Y\right)$
and the subspaces $\{gH_{j}\setminus Y\}_{g\in R(n-1)}$ are $L$-separated.
Here we recall the notion of decomposition of metric spaces from [GTY12] and [GTY13].
Definition 4.3.
An $r$-decomposition of a metric space $X$ over a metric family $\mathcal{Y}$ is a decomposition
$X=X_{0}\cup X_{1},\quad X_{i}=\bigsqcup_{j}X_{i,j}$
where each $X_{i,j}\in\mathcal{Y}$ and the distance between distinct $X_{i,j}$ and $X_{i,j^{\prime}}$ is grater than $r$ for $i=0,1$. A metric family $\mathcal{X}$ is $r$-decomposable over $\mathcal{Y}$ if each $X\in\mathcal{X}$ admits an $r$-decomposition over $\mathcal{Y}$.
Definition 4.4.
Let $\mathfrak{A}$ be a collection of metric families. A metric family $\mathcal{X}$ is decomposable over $\mathfrak{A}$, if for every $r>0$ there exists $\mathcal{Y}\in\mathfrak{A}$ such that $\mathcal{X}$ admits an $r$-decomposition over $\mathcal{Y}$.
Examples 4.5 (Theorem 4.1 of [GTY13]).
Let $\mathfrak{C}$ be the smallest collection of metric families, which contains all uniformly bounded metric families and closed under decomposition (i.e. if $\mathcal{Y}$ decomposes over $\mathfrak{C}$, then $\mathcal{Y}\in\mathfrak{C}$). A metric space $X$ is said to have finite decomposition complexity if the metric family $\{X\}$ is in $\mathfrak{C}$.
If a metric space $X$ has finite asymptotic dimension (not necessarily with bounded geometry), then $X$ has finite decomposition complexity. Especially, the metric family $\{(G,d_{s\cup\bigcup H_{i}})\}$ has finite decomposition complexity.
Let us clarify what we need to prove to show Theorem 1.3. For this we define a class of metric family.
Definition 4.6.
A metric family $\mathcal{X}$ is said to satisfy the twisted coarse Baum–Connes conjecture uniformly with respect to any stable coarse algebras if for any stable coarse $\bigsqcup_{X\in\mathcal{X}}X$-algebra $\Gamma(\bigsqcup_{X\in\mathcal{X}}X,\mathcal{A}\otimes\mathcal{K})$, we have
[…]
We denote by $\mathfrak{D}$ the class of metric families $\mathcal{X}$ which consists of subspaces of $G$ which satisfies the twisted coarse Baum–Connes conjecture uniformly with respect to any stable coarse algebras.
Since $(G,d_{s\cup\bigcup H_{i}})$ has finite decomposition complexity, to prove that $\{(G,d_{S})\}\in\mathfrak{D}$, it suffices to show the following lemmas:
Lemma 4.7.
Assume each peripheral subgroup $H_{i}$ satisfies the twisted coarse Baum–Connes conjecture with respect to any stable coarse algebras. If $\mathcal{X}$ is a metric family consists of subspaces of $(G,d_{S})$ whose diameter with respect to the distance $d_{s\cup\bigcup H_{i}}$ is uniformly bounded, then $\mathcal{X}\in\mathfrak{D}$.
Lemma 4.8.
If $\mathcal{X}$ decomposes over $\mathfrak{D}$, then $\mathcal{X}$ is in $\mathfrak{D}$.
We prove Lemma 4.7 in Section 5.3 and Lemma 4.8 in Section 5.2.
5 Proof of Lemma 4.7 and 4.8
In this section, we prove Lemma 4.7 and Lemma 4.8 to complete the proof of Theorem 1.3. The lemmas are consequences of controlled Mayer-Vietoris sequences, which first appeared in [Yu98]. The proof is divided into three parts. In the first part, we show that a decomposition of metric spaces induces a decomposition of $C^{*}$-algebras so that we can use the controlled Mayer-Vietoris sequence formulated in [GWY24]. Second, using the cutting-and-pasting argument, we reduce the problem to the twisted coarse Baum–Connes conjecture for $B(n)$ for each $n$. Finally, we show that any metric family consisting of subspaces of $B(n)$ satisfies the twisted coarse Baum–Connes conjecture with respect to any stable coarse algebras again using the cutting-and-pasting argument.
5.1 Decomposition
Let $(X,d)$ be a metric space and assume that $Y_{0},Y_{1}\subset X$ satisfies $X=Y_{0}\cup Y_{1}$. For each $r>0$ and $i=0,1$, we consider the equivalence relation on $Y_{i}$ generated by
$x\sim y\Leftrightarrow d(x,y)<3r$
and decompose $Y_{i}=\bigsqcup_{j}Y_{i}^{(j)}$ into $3r$-separated equivalence classes. On each disjoint union $\bigsqcup_{j}N_{r}(Y_{i}^{(j)})$, we consider the distance
$d_{i}(x,y)=\left\{\begin{array}[]{cc}d(x,y)&x\in N_{r}(Y_{i}^{(j)}),~y\in N_{r}(Y_{i}^{(j)})\text{ for some }j\\ \infty&\text{otherwise.}\end{array}\right$.
We denote by $Y_{i}^{+r}$ the space $\bigsqcup_{j}N_{r}(Y_{i}^{(j)})$ equipped with the distance $d_{i}$.
For any coarse $X$-algebra $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$ and a subspace $X^{\prime}\subset X$, we define a coarse $X^{\prime}$-algebra $\Gamma(X^{\prime},\mathcal{A}\otimes\mathcal{K})$ by
$\Gamma(X^{\prime},\mathcal{A}\otimes\mathcal{K}):=\{f|_{X^{\prime}}:f\in\Gamma(X,\mathcal{A}\otimes\mathcal{K})\}.$ (9)
Fix $s_{0}\geq 1$ and for each decomposition $\omega=(Y_{0},Y_{1})$ of $X$, define filtered subspaces by
$B_{r}^{\omega}$ $=C^{*}_{L,0}(P_{s_{0}}(Y_{0}^{+r});\Gamma(Y_{0}^{+r},\mathcal{A}\otimes\mathcal{K}))_{r}+C^{*}_{L,0}(P_{s_{0}}(Y_{1}^{+r});\Gamma(Y_{1}^{+r},\mathcal{A}\otimes\mathcal{K}))_{r}$
$I_{r}^{\omega}$ $=C^{*}_{L,0}(P_{s_{0}}(Y_{0}^{+r});\Gamma(Y_{0}^{+r},\mathcal{A}\otimes\mathcal{K}))_{r}$
$J_{r}^{\omega}$ $=C^{*}_{L,0}(P_{s_{0}}(Y_{1}^{+r});\Gamma(Y_{1}^{+r},\mathcal{A}\otimes\mathcal{K}))_{r}$
and define a $C^{*}$-subalgebra $B^{\omega}:=\overline{\bigcup_{r\geq 0}B_{r}^{\omega}}$ of $C^{*}_{L,0}(P_{s_{0}}(X);\Gamma(X,\mathcal{A}\otimes\mathcal{K}))$ and two ideals $I^{\omega}:=\overline{\bigcup_{r\geq 0}I_{r}^{\omega}(X)}$ and $J^{\omega}:=\overline{\bigcup_{r\geq 0}J_{r}^{\omega}(X)}$ of $B^{\omega}$.
Lemma 5.1.
The family of triples $(I^{\omega},J^{\omega};B^{\omega})_{\omega}$ is uniformly excisive (Definition 3.9), where $\omega$ runs all decompositions $\omega=(Y_{0},Y_{1})$ of $X$.
Proof.
We prove the first condition in Definition 3.9. Decompose $P_{s_{0}}(Y_{0}^{+r})\cup P_{s_{0}}(Y_{1}^{+r})$ into $Z_{0}:=P_{s_{0}}(Y_{0}^{+r})\setminus P_{s_{0}}(Y_{1}^{+r})$, $Z_{1}:=P_{s_{0}}(Y_{1}^{+r})\setminus P_{s_{0}}(Y_{0}^{+r})$ and $Z_{0,1}:=P_{s_{0}}(Y_{0}^{+r})\cap P_{s_{0}}(Y_{1}^{+r})$. For any $A\in B_{r}^{\omega}$, define
$A_{0}$ $:=\chi_{Z_{0}}A\chi_{Z_{0}}+\chi_{Z_{0}}A\chi_{Z_{0,1}}+\chi_{Z_{0,1}}A\chi_{Z_{0}}+\chi_{Z_{0,1}}A\chi_{Z_{0,1}}$
$A_{1}$ $:=\chi_{Z_{1}}A\chi_{Z_{1}}+\chi_{Z_{1}}A\chi_{Z_{0,1}}+\chi_{Z_{0,1}}A\chi_{Z_{1}}.$
Then $A=A_{0}+A_{1}$, $A_{0}$ and $A_{1}$ have propagation less than $r$, $A_{0}$ is supported on $P_{s_{0}}(Y_{0}^{+r})\times P_{s_{0}}(Y_{0}^{+r})$ and $A_{1}$ is supported on $P_{s_{0}}(Y_{1}^{+r})\times P_{s_{0}}(Y_{1}^{+r})$. Moreover we have
$\|A_{0}\|\leq 4\|A\|,~\|A_{1}\|\leq 3\|A\|.$
For any partial translation $v$ on $Y_{0}^{+r}$, the functions $(A_{0})^{v}$ is in $\Gamma(Y_{0}^{+r},\mathcal{A}\otimes\mathcal{K})$ since it is a restriction of $A^{v}$ by the definition of the coarse algebra (9). Therefore $A_{0}\in I_{r_{0}}^{\omega}$. Similarly, we have $A_{1}\in J_{r_{0}}^{\omega}$.
Next, we prove the second condition. For $r>0$, we define a partition of unity $\{\mu_{0}^{(r)},\mu_{1}^{(r)}\}$ of $P_{s_{0}}(Y_{0}^{+r})\cup P_{s_{0}}(Y_{1}^{+r})$ by
$\mu_{i}^{(r)}(x):=\frac{d\left(x,{P_{s_{0}}(Y_{i}^{+r})}^{C}\right)}{d\left(x,{P_{s_{0}}(Y_{0}^{+r})}^{C}\right)+d\left(x,{P_{s_{0}}(Y_{1}^{+r})}^{C}\right)}.$
Since
$d\left(x,{P_{s_{0}}(Y_{1}^{+r})}^{C}\right)+d\left(x,{P_{s_{0}}(Y_{2}^{+r})}^{C}\right)\geq r,$
and
$d(x,y)\leq r_{0}\Rightarrow\left|d\left(x,{P_{s_{0}}(Y_{i}^{+r})}^{C}\right)-d\left(y,{P_{s_{0}}(Y_{i}^{+r})}^{C}\right)\right|\leq r_{0},$
it is easy to see that
$d(x,y)\leq r_{0}\Rightarrow\left|\mu_{i}^{(r)}(x)-\mu_{i}^{(r)}(y)\right|\leq\frac{r_{0}}{r}.$
Let $a\in I^{\omega}\otimes\mathcal{K}\cap J^{\omega}\otimes\mathcal{K}$ such that $d(a,a_{1})<\delta$ and $d(a,a_{2})<\delta$ for some $a_{1}\in(I^{\omega}\otimes\mathcal{K})_{r_{0}}$ and $a_{2}\in(J^{\omega}\otimes\mathcal{K})_{r_{0}}$. Then,
$\left\|[\mu_{i}^{(r)},a_{j}]_{x,y}\right\|=\left\|(a_{j})_{x,y}\left(\mu_{i}^{(r)}(x)-\mu_{i}^{(r)}(y)\right)\right\|\leq\|a_{j}\|\cdot\left\|\mu_{i}^{(r)}(x)-\mu_{i}^{(r)}(y)\right\|.$
Fix any $\varepsilon>0$. By [Roe03, Lemma 4.27], there exists a constant $C(r_{0})$ such that
$\left\|[\mu_{i}^{(r)},a_{j}]\right\|\leq C(r_{0})\frac{r_{0}}{r}\|a_{j}\|<\frac{\varepsilon}{4}$
by taking $r>\frac{4r_{0}C(r_{0})}{\varepsilon}\|a_{j}\|$. Now define
$b:=\sqrt{\mu_{2}}a_{1}\sqrt{\mu_{2}}+\sqrt{\mu_{1}}a_{2}\sqrt{\mu_{1}}.$
Then $b$ is supported on $\left(P_{s_{0}}(Y_{0}^{+r})\times P_{s_{0}}(Y_{0}^{+r})\right)\cap\left(P_{s_{0}}(Y_{1}^{+r})\times P_{s_{0}}(Y_{1}^{r})\right)$. Therefore, by the third condition of Definition 2.1, we have
$b\in I^{\omega}_{r}\otimes\mathcal{K}\cap J^{\omega}_{r}\otimes\mathcal{K}.$
In addition, if $\delta<\frac{\varepsilon}{4}$, we have
$\|a-b\|$
$\leq$ $\|a-\left(\sqrt{\mu_{2}}a\sqrt{\mu_{2}}+\sqrt{\mu_{1}}a\sqrt{\mu_{1}}\right)\|+\|(\sqrt{\mu_{2}}a\sqrt{\mu_{2}}+\sqrt{\mu_{1}}a\sqrt{\mu_{1}})-(\sqrt{\mu_{2}}a_{1}\sqrt{\mu_{2}}+\sqrt{\mu_{1}}a_{2}\sqrt{\mu_{1}})\|$
$\leq$ $\frac{\varepsilon}{2}+2\delta<\varepsilon$
Therefore, the family of triples $(B^{\omega};I^{\omega},J^{\omega})_{\omega}$ is uniformly excisive. ∎
Moreover, for $s\geq s_{0}$, define
$B_{r}^{\omega,s}$ $=C^{*}_{L,0}(P_{s_{0}}(Y_{0}^{+r});\Gamma(Y_{0}^{+r},\mathcal{A}\otimes\mathcal{K}))_{r}+C^{*}_{L,0}(P_{s_{0}}(Y_{1}^{+r});\Gamma(Y_{1}^{+r},\mathcal{A}\otimes\mathcal{K}))_{r}$
$\quad\quad\quad\quad+C^{*}_{L,0}(P_{s_{0}}(Y_{0}^{+r})\cap P_{s_{0}}(Y_{1}^{+r}));\Gamma(Y_{0}^{+r}\cap Y_{1}^{+r},\mathcal{A}\otimes\mathcal{K}))_{sr}$
$I_{r}^{\omega,s}$ $=C^{*}_{L,0}(P_{s_{0}}(Y_{0}^{+r});\Gamma(Y_{0}^{+r},\mathcal{A}\otimes\mathcal{K}))_{r}+C^{*}_{L,0}(P_{s}(Y_{0}^{+r}\cap Y_{1}^{+r});\Gamma(Y_{0}^{+r}\cap Y_{1}^{+r},\mathcal{A}\otimes\mathcal{K}))_{sr}$
$J_{r}^{\omega,s}$ $=C^{*}_{L,0}(P_{s_{0}}(Y_{1}^{+r});\Gamma(Y_{1}^{+r},\mathcal{A}\otimes\mathcal{K}))_{r}+C^{*}_{L,0}(P_{s}(Y_{0}^{+r}\cap Y_{1}^{+r});\Gamma(Y_{0}^{+r}\cap Y_{1}^{+r},\mathcal{A}\otimes\mathcal{K}))_{sr}$
define a $C^{*}$-subalgebra $B^{\omega,s}:=\overline{\bigcup_{r\geq 0}B_{r}^{\omega,s}}$ of $C^{*}_{L,0}(P_{s_{0}};\Gamma(X,\mathcal{A}\otimes\mathcal{K}))$ and two ideals $I^{\omega,s}:=\overline{\bigcup_{r\geq 0}I_{r}^{\omega,s}}$ and $J^{\omega,s}:=\overline{\bigcup_{r\geq 0}J_{r}^{\omega,s}}$ of $B^{\omega,s}$.
Lemma 5.2.
The family of triples $(I^{\omega,s}(X),J^{\omega,s}(X);B^{\omega,s}(X))_{\omega,s}$ is uniformly excisive, where $\omega$ runs all decompositions $\omega=(Y_{1},Y_{2})$ of $X$ and $s$ runs all $s\geq s_{0}$.
Proof.
Same as the proof of Lemma 5.1. ∎
We have a preliminary lemma (analogous to [GWY24, Lemma 7.14]).
Lemma 5.3.
Let $X$ be a metric space and $r_{0},s_{0}>0$. Assume that $X$ admits a decomposition $X=Y_{0}\cup Y_{1}$. Then we have
$C^{*}_{L,0}(P_{s_{0}}(X);\Gamma(X,\mathcal{A}\otimes\mathcal{K}))_{r_{0}}\subset C^{*}_{L,0}$ $\left(P_{s_{0}}(Y_{0}^{+(2s_{0}+r_{0})});\Gamma(Y_{0}^{+(2s_{0}+r_{0})},\mathcal{A}\otimes\mathcal{K})\right)_{r_{0}}$
$\quad\quad+C^{*}_{L,0}\left(P_{s_{0}}(Y_{1}^{+(2s_{0}+r_{0})}),\Gamma(Y_{1}^{+(2s_{0}+r_{0})};\mathcal{A}\otimes\mathcal{K})\right)_{r_{0}}.$
Proof.
We have
$P_{s_{0}}(X)\subset P_{s_{0}}\left(Y_{0}^{+s_{0}}\right)\cup P_{s_{0}}\left(Y_{1}^{+s_{0}}\right).$
We denote by $\chi_{0}$ the characteristic function of $N_{r_{0}}(P_{s_{0}}(Y_{0}^{+s_{0}}))\subset P_{s_{0}}(X)$. For any $A\in C^{*}_{L,0}(P_{s_{0}}(X))_{r_{0}}$, we define
$A_{0}=\chi_{0}A\chi_{0},~A_{1}=A-A_{0}.$
Then $A_{0}$ is supported on $N_{r_{0}}(P_{s_{0}}(Y_{0}^{+s_{0}}))\times N_{r_{0}}(P_{s_{0}}(Y_{0}^{+s_{0}}))$ and
$A_{1}=A(1-\chi_{0})+(1-\chi_{0})A\chi_{0}$
is supported on $N_{r_{0}}(P_{s_{0}}(Y_{1}^{+s_{0}}))\times N_{r_{0}}(P_{s_{0}}(Y_{1}^{+s_{0}}))$ since the propagation of $A$ is less than $r_{0}$. So we have
$C^{*}_{L,0}(P_{s_{0}}(X))_{r_{0}}\subset C^{*}_{L,0}\left(N_{r_{0}}\left(P_{s_{0}}\left(Y_{0}^{+s_{0}}\right)\right)\right)_{r_{0}}+C^{*}_{L,0}\left(N_{r_{0}}\left(P_{s_{0}}\left(Y_{1}^{+s_{0}}\right)\right)\right)_{r_{0}}.$
But for $i=0,1$, we have
$N_{r_{0}}\left(P_{s_{0}}\left(Y_{i}^{+s_{0}}\right)\right)\subset P_{s_{0}}\left(Y_{i}^{+(2s_{0}+r_{0})}\right).$
Therefore we have
$C^{*}_{L,0}(P_{s_{0}}(X))_{r_{0}}\subset C^{*}_{L,0}\left(P_{s_{0}}\left(Y_{0}^{+(2s_{0}+r_{0})}\right)\right)_{r_{0}}+C^{*}_{L,0}\left(P_{s_{0}}\left(Y_{1}^{+(2s_{0}+r_{0})}\right)\right)_{r_{0}}.$
As well as the first half of the proof of Lemma 5.1, if $A\in C^{*}_{L,0}(P_{s_{0}}(X);\Gamma(X,\mathcal{A}\otimes\mathcal{K}))_{r_{0}}$, we have
$A_{i}\in C^{*}_{L,0}\left(P_{s_{0}}(Y_{i}^{+(2s_{0}+r_{0})});\Gamma(Y_{0}^{+(2s_{0}+r_{0})},\mathcal{A}\otimes\mathcal{K})\right)_{r_{0}}$
for $i=0,1$ again since they are restrictions. ∎
5.2 Reduction to $(B(n),d_{S})$
In this section, we prove Lemma 4.8. Let $x\in K_{*}^{r_{0},\frac{1}{8}}(\mathbb{C}_{L,0}(P_{s_{0}}(\bigsqcup X);\Gamma(\bigsqcup X,\mathcal{A}\otimes\mathcal{K}))$ be any element for any $r_{0}$ and $s_{0}$, where $\Gamma\left(\bigsqcup X,\mathcal{A}\otimes\mathcal{K}\right)$ is any stable coarse $\left(\bigsqcup_{X\in\mathcal{X}}X\right)$-algebra. It suffices to show that there exists $s\geq s_{0}$ such that $x=0$ in
$K_{*}^{s,\frac{1}{8}}\left(C^{*}_{L,0}\left(P_{s}\left(\bigsqcup X\right);\Gamma\left(\bigsqcup X,\mathcal{A}\otimes\mathcal{K}\right)\right)\right).$
For any subset $X^{\prime}\subset\bigsqcup_{X\in\mathcal{X}}X$, $\Gamma(X^{\prime},\mathcal{A}\otimes\mathcal{K})$ is the coarse $X^{\prime}$-algebra defined in (9). Decompose
$X=\bigsqcup_{j}Y_{0}^{(j)}(X)\cup\bigsqcup_{j}Y_{1}^{(j)}(X)$
such that the family $\{Y_{i}^{(j)}(X)\}_{j}$ is $3(2r_{0}+s_{0})$-separated and each $Y_{i}^{(j)}(X)$ is in $\mathcal{Y}$ for $i=0,1$ for some $\mathcal{Y}\in\mathfrak{D}$. Then this gives a decomposition
$\bigsqcup_{X\in\mathcal{X}}X=\bigsqcup_{j,X}Y_{0}^{(j)}(X)\cup\bigsqcup_{j,X}Y_{1}^{(j)}(X).$
We denote this decomposition of $\bigsqcup_{X\in\mathcal{X}}X$ by $\omega=\left(\bigsqcup_{j,X}Y_{0}^{(j)}(X),\bigsqcup_{j,X}Y_{1}^{(j)}(X)\right)$ and denote $Y_{i}:=\bigsqcup_{j,X}Y_{i}^{(j)}(X)$ where the distance between distinct components is taken to be infinite. By Lemma 5.3, $C^{*}_{L,0}\left(P_{s_{0}}(\bigsqcup X);\Gamma(\bigsqcup X,\mathcal{A}\otimes\mathcal{K})\right)_{r_{0}}$ is contained in
$C^{*}_{L,0}\left(P_{s_{0}}\left(Y_{0}^{+(2s_{0}+r_{0})}\right);\Gamma\left(Y_{0}^{+(2s_{0}+r_{0})},\mathcal{A}\otimes\mathcal{K}\right)\right)_{r_{0}}+C^{*}_{L,0}\left(P_{s_{0}}\left(Y_{0}^{+(2s_{0}+r_{0})}\right);\Gamma\left(Y_{0}^{+(2s_{0}+r_{0})},\mathcal{A}\otimes\mathcal{K}\right)\right)_{r_{0}},$
which is contained in $B_{2s_{0}+r_{0}}^{\omega}$. Therefore, $x$ is sent to an element in $K^{\frac{1}{8},2s_{0}+r_{0}}(B^{\omega})$. There exists $r_{1}\geq 2s_{0}+r_{0}$ and the boundary map $\partial$, which fits in the commutative diagram below for every $s_{1}\geq s_{0}$:
$K_{*}^{2r_{0}+s_{0},\frac{1}{8}}(B^{\omega})$$K_{*}^{r_{1},\frac{1}{8}}(I^{\omega}\cap J^{\omega})$$K_{*}^{2r_{0}+s_{0},\frac{1}{8}}(B^{\omega,s_{1}})$$K_{*}^{r_{1},\frac{1}{8}}(I^{\omega,s_{1}}\cap J^{\omega,s_{1}}).$$\scriptstyle{\partial}\scriptstyle{\partial}$
Since $\Gamma(Y_{0}^{+r},\mathcal{A}\otimes\mathcal{K})\cap\Gamma(Y_{1}^{+r},\mathcal{A}\otimes\mathcal{K})=\Gamma(Y_{0}^{+r}\cap Y_{1}^{+r},\mathcal{A}\otimes\mathcal{K})$, the right vertical map can be identified with
$K_{*}^{r_{1},\frac{1}{8}}(C^{*}_{L,0}$ $(P_{s_{0}}(Y_{0}^{+r_{1}}\cap Y_{1}^{+r_{1}});\Gamma(Y_{0}^{+r_{1}}\cap Y_{1}^{+r_{1}},\mathcal{A}\otimes\mathcal{K}))$
$\longrightarrow K_{*}^{r_{1}s_{1},\frac{1}{8}}(C^{*}_{L,0}(P_{s_{1}}(Y_{0}^{+r_{1}}\cap Y_{1}^{+r_{1}});\Gamma(Y_{0}^{+r_{1}}\cap Y_{1}^{+r_{1}},\mathcal{A}\otimes\mathcal{K}))).$
By the coarse invariance and the permanence under taking subspace of the twisted coarse Baum–Connes conjecture with respect to any stable coarse algebras, $Y_{0}^{+r_{1}}\cap Y_{1}^{+r_{1}}$ satisfies the twisted coarse Baum–Connes conjecture with respect to any stable coarse algebras. Therefore, $\partial x$ is sent to 0 for sufficiently large $s_{1}$. By the controlled-exactness of the Mayer-Vietoris sequence, there exists $r_{2}$ and
$y\in K_{*}^{r_{2},\frac{1}{8}}(I^{\omega,s_{1}}),~z\in K_{*}^{r_{2},\frac{1}{8}}(J^{\omega,s_{1}})$
such that $x=y+z$ in $K_{*}^{r_{2},\frac{1}{8}}(B^{\omega,s_{1}})$. Since
$I^{\omega,s_{1}}_{r_{2}}\subset I^{\omega}_{s_{1}r_{2}},$
by the inductive assumption, $y=0$ in $K_{*}^{s,\frac{1}{8}}(I^{\omega})$ for sufficiently large $s$ and the same thing holds for $z$. Therefore, we have
$x=y+z=0\in K_{*}^{s,\frac{1}{8}}\left(C^{*}_{L,0}\left(P_{s}\left(\bigsqcup X\right);\Gamma\left(\bigsqcup X;\mathcal{A}\otimes\mathcal{K})\right)\right)\right).$
5.3 The twisted coarse Baum–Connes conjecture for $B(n)$
In this subsection, we prove Lemma 4.7.
Lemma 5.4.
Let $\{Y_{j}\}_{j=1}^{\infty}$ be a sequence of subspaces of $X$. If $X$ satisfies the twisted coarse Baum–Connes conjecture with respect to any stable coarse algebras, then $\{Y_{j}\}$ satisfies the twisted coarse Baum–Connes conjecture uniformly with respect to any stable coarse algebras.
Proof.
This follows from Corollary 2.24 and Lemma 2.28. ∎
We record a useful corollary, which is very specific to the twisted coarse Baum–Connes conjecture.
Corollary 5.5.
Let $X\subset Y_{1},Y_{2}$. If $Y_{1}$ and $Y_{2}$ satisfy the twisted coarse Baum–Connes conjecture with respect to any stable coarse algebras, then so does $Y_{1}\cup Y_{2}$.
Proof.
This follows from Lemma 4.8. ∎
Corollary 5.6.
Assume $G$ is hyperbolic relative to $\{H_{i}\}_{i=1}^{N}$. If every $H_{i}$ satisfies the twisted coarse Baum–Connes conjecture with any stable coarse algebras, then so does $B(n)$.
Proof.
Induction on $n$. When $n=1$, use Lemma 5.4 and Corollary 5.5 to
$B(1)=S\cup\left(\bigcup_{i}H_{i}\right).$
Assume the conclusion holds for $1,2,\cdots,n-1$. Denote
$B(n;i)=B(n-1)\cdot S\cup B(n-1)\cdot H_{i}.$
Since $B(n)=\bigcup_{i=1}^{n}B(n;i)$, it suffices to show that $B(n;i)$ satisfies the twisted coarse Baum–Connes conjecture with any stable coarse algebras, again by Corollary 5.5. But this follows from Lemma 4.2, Lemma 4.8 and Corollary 5.5. ∎
Now Lemma 4.7 follows from Lemma 5.4 and Corollary 5.6.
Acknowledgements
We would like to thank Prof. Guoliang Yu for his comments and discussions on this topic.